### v5

# New Permutation Decomposition Techniques For Efficient Homomorphic Permutation

Abstract.
Homomorphic permutation is fundamental to privacy-preserving computations based on batch-encoding homomorphic encryption. It underpins nearly all homomorphic matrix operations and predominantly influences their complexity. Permutation decomposition as a potential approach to optimize this critical component remains underexplored. In this paper, we propose novel decomposition techniques to optimize homomorphic permutations, advancing homomorphic encryption-based privacy-preserving computations.
We start by defining an ideal decomposition form for permutations and propose an algorithm searching depth-1 ideal decompositions. Based on this, we prove the full-depth ideal decomposability of permutations used in specific homomorphic matrix transposition (HMT) and multiplication (HMM) algorithms, allowing them to achieve asymptotic improvement in speed and rotation key reduction. As a demonstration of applicability, substituting the HMM components in the best-known inference framework of encrypted neural networks with our enhanced version shows up to 7.9×7.9$\times$7.9 × reduction in latency.
We further devise a new method for computing arbitrary homomorphic permutations, specifically those with weak structures that cannot be ideally decomposed. We design a network structure that deviates from the conventional scope of decomposition and outperforms the state-of-the-art technique with a speed-up of up to 1.69×1.69$\times$1.69 × under a minimal rotation key requirement.
Homomorphic Encryption, Homomorphic Permutation, Permutation Decomposition
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Cryptography
1. Introduction
Homomorphic encryption (HE) is a promising cryptographic primitive that facilitates computations on encrypted data without revealing the underlying plaintext. Significant technical advances in HE have been made following Gentry’s breakthrough (Gentry, 2009). Among these is a class of HE schemes supporting SIMD (single instruction, multiple data) techniques (Gentry et al., 2012; Smart and Vercauteren, 2014). These schemes (Brakerski and Vaikuntanathan, 2011; Fan and Vercauteren, 2012; Brakerski et al., 2014; Halevi et al., 2019; Cheon et al., 2017) allow multiple plaintext values to be packed as a vector(or matrix) and encrypted into a single ciphertext, enabling homomorphic coordinate-wise additions, multiplications, and collective shifts (known as rotations) of vector components. The advantages of batch encoding and parallel computation result in good amortized complexity per plaintext value, making them applicable in various privacy-preserving computation scenarios.
In HE-based privacy-preserving computations, particularly those related to machine learning, matrix operations on encrypted data are required as core computing components. For instance, homomorphic matrix multiplication and transposition are indispensable for the homomorphic evaluation of neural network propagation algorithms (Gilad-Bachrach et al., 2016; Jiang et al., 2018; Lee et al., 2022; Sav et al., 2021, 2022; Brutzkus et al., 2019).
In the plaintext computational scenario, matrix operations need to retrieve and arrange matrix entries frequently by index addressing. However, for matrices encrypted under a batch-encoding HE scheme, this task can only be achieved through basic ciphertext operations like addition, multiplication, and rotation. We refer to this process as homomorphic permutation.
Homomorphic permutation often dominates the complexity of homomorphic matrix operations, making its optimization critical for improving the performance of HE-based privacy-preserving applications. Specifically, homomorphic permutation necessitates a substantial number of ciphertext rotations (Halevi and Shoup, 2014, 2018) which are among the most computationally intensive operations in HE, involving multiple number-theoretic transformations over the polynomial ring. Moreover, distinct rotation keys are required to achieve rotations of varying distances, each of which can be significantly larger than the ciphertext itself. The need for numerous rotation keys introduces considerable overhead in both communication and storage (Lee et al., 2023; Chan et al., 2024).
One common optimization for homomorphic permutation involves deconstructing ciphertext rotations and reordering subroutine sequences to balance time complexity with the number of rotation keys. This includes the hoisting technique proposed by Halevi and Shoup (Halevi and Shoup, 2018) and its extended version known as double hoisting for homomorphic linear transformation (Bossuat et al., 2021).
Another compatible optimization is permutation decomposition. It aims to decompose a permutation into multiple simpler ones, each requiring fewer rotations and keys. This concept was first proposed by Gentry et al. (Gentry et al., 2012) where they used the Benes network (Beneš, 1964) for general permutation decomposition. Yet, the application of permutation decomposition in privacy-preserving computations is limited (Meftah et al., 2021, 2022), as existing general methods (Gentry et al., 2012; Halevi and Shoup, 2014) lack adaptability for practical permutations.
In this paper, we introduce new techniques to devise more effective decompositions, achieving enhancement in both specific and general use cases of homomorphic permutation.
1.1. Technique Overview & Contributions
We begin by analyzing the potential optimal enhancements that decompositions can bring to homomorphic permutations. This leads us to define the ideal decomposition form, which adaptively reduces the complexity of the homomorphic permutation from polynomial to logarithmic, based on the non-zero diagonal distribution of the permutation matrix (Section 2). The rest of the paper revolves around achieving or closely approximating the defined ideal optimization for specific and arbitrary permutations:
Firstly, we present an algorithm that searches depth-1 ideal decomposition solutions for arbitrary permutations (Section 3). This enables us to discover the full-depth ideal decomposability of specific permutations in homomorphic matrix transposition (HMT) proposed by Jiang et al.(Jiang et al., 2018) and homomorphic matrix multiplication (HMM) by Rizomiliotis and Triakosia (Rizomiliotis and Triakosia, 2022) (Section 4), allowing these privacy-preserving computations to achieve asymptotic improvement in speed and rotation key reduction. We also present the application of our enhanced HMM, which performs inference on an encrypted convolutional neural network model. By substituting the HMM component in the state-of-the-art solution with our version, we achieve a significant improvement in latency. More usage of the search algorithm for decomposing other valuable permutations are provided in the appendix (Appendix C).
Secondly, since not every permutation guarantees ideal decomposability, we further devise a new algorithm for decomposing arbitrary homomorphic permutations (Section 5). This aims to better approximate the performance of ideal decomposition than previous approaches. We construct a multi-group network structure with nodes taking vectors as input and output, applying rotations with varying steps. Through precise node transmission logic and tailored homomorphic operations, our approach outperforms the state-of-the-art implementation using the Benes network (Gentry et al., 2012; Halevi and Shoup, 2014) under minimal rotation key requirements.
2. Background and Preliminary
2.1. Notations
We denote scalars and polynomials by lowercase letters, vectors by bold lowercase letters, and matrices by uppercase letters. $A[i,j]$ represents the element in the $i$-th row and $j$-th column of matrix $A$. We also use $A(k,l)_{r}$ to indicate the element in the $k$-th diagonal and $l$-th row of $A$, where $A(k,l)_{r}=A[l,k+i]$. Note that index computation is always implicitly reduced by the matrix’s dimension. Moreover, we denote element-wise product by $\odot$.
2.2. Batch-encoding Homomorphic Encryption
As mentioned above, the batch-encoding homomorphic encryption is a class of HE schemes (Brakerski and Vaikuntanathan, 2011; Gentry et al., 2012; Fan and Vercauteren, 2012; Cheon et al., 2017) based on the security assumption of RLWE (Lyubashevsky et al., 2010, 2013), which supports vector spaces as plaintext spaces and enables homomorphic element-wise operations.
Typically, for a vector $\mathbf{z}$ to be encrypted, a batch-encoding HE scheme first maps it to some polynomial ring space $R=\mathbb{Z}[X]/(X^{N}+1)$ with $N$ a power of $2$. The mapped plaintext polynomial, denoted $pt(\mathbf{z})$, is then encrypted into a ciphertext $ct(\mathbf{z})$ in $R_{Q_{\ell}}^{2}=(\mathbb{Z}_{Q_{\ell}}[X]/(X^{N}+1))^{2}$ where $Q_{\ell}=\prod_{i=0}^{\ell}q_{i},0\leq\ell<L$ for some pre-determined modulus chain $Q=[q_{0},\dots,q_{L-1}]$. Such ciphertext generally supports the following homomorphic operations (see Appendix A for more details):
  • $\bullet$
    $\texttt{Add}(ct(\mathbf{z}_{1}),ct(\mathbf{z}_{2}))$: For the input ciphertexts $ct(\mathbf{z}_{1})\in R_{Q_{\ell}}^{2}$ and $ct(\mathbf{z}_{2})\in R_{Q_{\ell^{\prime}}}^{2}$, output $ct(\mathbf{z}_{1})+ct(\mathbf{z}_{2})\bmod{Q_{\min(\ell,\ell^{\prime})}}$, which is a valid ciphertext of $\mathbf{z}_{1}+\mathbf{z}_{2}$.
  • $\bullet$
    $\texttt{CMult}(ct(\mathbf{z}_{1}),pt(\mathbf{z}_{2}))$: For the input ciphertext $ct(\mathbf{z}_{1})\in R_{Q_{\ell}}^{2}$ and plaintext $pt(\mathbf{z}_{2})\in R_{Q_{\ell^{\prime}}}$, output $ct(\mathbf{z}_{1})\cdot pt(\mathbf{z}_{2})\bmod{Q_{\min(\ell,\ell^{\prime})}}$ as a valid ciphertext of $\mathbf{z}_{1}\odot\mathbf{z}_{2}$.
  • $\bullet$
    $\texttt{Mult}(ct(\mathbf{z}_{1}),ct(\mathbf{z}_{2}),rlk)$: For the input ciphertexts $ct(\mathbf{z}_{1})\in R_{Q_{\ell}}^{2}$ and $ct(\mathbf{z}_{2})\in R_{Q_{\ell^{\prime}}}^{2}$, compute the tensor product $ct(\mathbf{z}_{1})\otimes ct(\mathbf{z}_{2})\bmod{Q_{\min(\ell,\ell^{\prime})}}$ in $R_{Q_{\min(\ell,\ell^{\prime})}}^{3}$. Then switch it back to $R_{Q_{\min(\ell,\ell^{\prime})}}^{2}$ using a relinearization key $rlk$ and output it as a valid ciphertext of $\mathbf{z}_{1}\odot\mathbf{z}_{2}$.
  • $\bullet$
    $\texttt{Rescale}(ct)$: For the input ciphertext $ct\in R_{Q_{\ell}}^{2}$, Rescale reduces the scaling factor of the underlying plaintext by $\frac{Q_{\ell}-1}{Q_{\ell}}$, resulting in the output: $\lfloor\frac{Q_{\ell}-1}{Q_{\ell}}\cdot ct\rceil\bmod{Q_{\ell-1}}\in R_{Q_{\ell-1}}^{2}$. This operation is required after a Mult or CMult for noise control.
  • $\bullet$
    $\texttt{Rot}(ct(\mathbf{z}),k)$: Homomorphically left-shifting each component of $\mathbf{z}$ by $k\bmod{n}$ positions with $n$ being the dimension of $\mathbf{z}$. A rotation key $rtk_{k}$ is required as an implicit input.
We summarize some common constraints of batch-encoding homomorphic encryption schemes:
  1. (1)
    Batch-encoding HE only natively supports homomorphic operations with a limited multiplication depth of $L-1$, as $Q_{L}$ provides at most $L-1$ moduli for Rescale operations.
  2. (2)
[…]
  3. (3)
    A rotation key consists of multiple ciphertexts encrypting slices of some Galois automorphism of $s$ scaled by some gadget vector and modulus chain $P$ (Brakerski et al., 2014). A Large number of rotation keys may cause a great burden on key generation and a high communication overhead for transmitting the keys (Lee et al., 2023; Chan et al., 2024). This can be a significant efficiency factor in some multi-party computation scenarios (Aloufi et al., 2021; Mouchet, 2023; Kim et al., 2023).
2.3. Homomorphic Permutation
Given a ciphertext $ct(\mathbf{x})$ encrypting an $n$-dimensional plaintext vector $\mathbf{x}$ and a length-$n$ permutation $p$ to be applied to $\mathbf{x}$, the homomorphic permutation $ct(p(\mathbf{x}))$ can be represented as a homomorphic linear transformation (HLT) $ct(U\mathbf{x})$, where $U$ is the matrix representation of $p$. $ct(U\mathbf{x})$ can be computed by:
(1) $ct(U\mathbf{x})\leftarrow\sum_{i=0}^{n-1}\mathbf{u}_{i}\odot\texttt{Rot}(ct(\mathbf{x});i),$
where $\mathbf{u}_{i}$ denotes the $i$-th diagonal of $U$ (Halevi and Shoup, 2014). The BSGS (Baby-Step Giant-Step) algorithm reformulates this expression to reduce its complexity from $O(n)$ ciphertext rotations to $O(n_{1}+n_{2})$ with $n_{1}n_{2}=n$ as follows:
(2) $\begin{split}\sum_{i=0}^{n_{2}-1}\texttt{Rot}\left(\sum_{j=0}^{n_{1}-1}\left(\mathbf{u}_{i,j}\odot\texttt{Rot}(ct(\mathbf{x});j)\right);n_{1}\cdot i\right).\end{split}$
where $\mathbf{u}_{i,j}=\texttt{Rot}(\mathbf{u}_{n_{1}\cdot i+j};-n_{1}\cdot i)$. If there exists a common difference $a$ and range $d\leq\lfloor n/a\rfloor$ such that non-zero diagonals are distributed over an arithmetic sequence $s=\{a\cdot i\mid 0\leq i\leq d\}$, a slight modification of BSGS yields a rotation complexity of $O(d_{1}+d_{2})$ with $d_{1}d_{2}=d$. The hoisting techniques (Halevi and Shoup, 2018; Bossuat et al., 2021) provide the BSGS algorithm with a trade-off between speed and the number of rotation keys required. Broadly, they require $O(d_{1}+d_{2})$ rotation keys and enable the speed of HLT to increase with the BSGS ratio $d_{1}/d_{2}$ within a certain range (potentially $1\sim 16$).
2.3.1. Permutation Decomposition
Equations 1 and 2 imply that the complexity of homomorphic permutation is closely related to the number of non-zero diagonals in its matrix representation. As mentioned in Section 1, permutation decomposition aims to factorize the permutation into factors possessing sparser non-zero diagonals. This contains two potential benefits: reduced ciphertext rotations and rotation keys.
Gentry et al. (Gentry et al., 2012) utilized the Benes network for permutation decomposition in homomorphic permutation. For any permutation $U$ of length $n$, this approach executes $\log{n}$ rounds of decomposition. In the initial iteration, $U$ is factorized into three components: $U=U^{\sigma}U^{\rho}U^{\tau}$, with $U^{\rho}$ serving as the input of the next round. $U^{\sigma}$ and $U^{\tau}$ contain only non-zero diagonals in {±$\{\pm 2^{\log{n}-i-1},0\}$, where $i$ indicates the iteration round. To our knowledge, the Benes network-based technique (Gentry et al., 2012; Halevi and Shoup, 2014) remains the first and optimal general decomposition method for homomorphic permutation.
Although the Benes network reduces the rotation complexity to roughly $O(4\log{n})$, it fails to adjust complexity based on the distribution of non-zero diagonals, making it counterproductive for highly structured permutations. Another limitation of the Benes network is its significant depth overhead, which can easily exceed the depth supported by homomorphic encryption schemes. While the depth can be reduced by collapsing network levels (Halevi and Shoup, 2014), the complexity of the merged permutations increases as well.
Theorem 1.
(Ma et al., 2024) Let the $n\times n$ matrix $U$ be the result of multiplying square matrices $U_{L}$ and $U_{R}$ of the same size. Then, $U[i,j]=\sum_{a=0}^{n-1}{{U_{L}[i,a]\cdot U_{R}[a,j]}}$. Denote the diagonal index of $U[i,j]$ as $k=j-i$, the diagonal index of $U_{L}[i,a]$ as $k_{1}=a-i$, and the diagonal index of $U_{R}[a,j]$ as $k_{2}=j-a$. It follows that $k_{1}+k_{2}=k\bmod{n}$.
Subsequent work designing decompositions for specific permutations gives us new insights (Ma et al., 2024). It proposed two-factor decompositions in the form $U=U_{L}U_{R}$, which includes an intriguing pattern (Theorem 1). This implies that two-factor decompositions can be seen as performing subtraction on the diagonal indices. If $U$ has all its non-zero diagonals distributed within the interval $[-r,r]\mod{n}$, and we designate its factor, $U_{R}$, to possess non-zero diagonals exclusively at $\{r_{c},-r_{c},0\}$ ($r_{c}\leq r$), then $U_{L}$ may have its non-zero diagonals confined to the range $[-r^{\prime},r^{\prime}]$ where $r^{\prime}=\max{(r-r_{c},r_{c}/2)}$. Then $U$ can be substituted by $U_{L}$ for further decomposition, as $U_{R}$ maintains a constant sparsity. This idea allows $U$ to be recursively decomposed into a chain:
(3) $U=U_{L_{\ell}}U_{R_{\ell}}U_{R_{\ell-1}}\dots U_{R_{1}}.$
The non-zero diagonals of $U_{L_{\ell}}$ progressively contract from both sides of the 00-th diagonal, becoming sparser as $\ell$ increases.
In this paper, we specifically define an ideal decomposition form (Definition 2). Its $O(2\log(r/a))$ rotations and $O(\log(r/a))$ depth indicate that $n$ has been decoupled from the complexity, highlighting an enhanced adaptability to highly structured permutations. Such an optimization effect may represent a potential upper bound achievable through decomposition techniques.
Definition 0.
Given a square matrix $U$ with non-zero diagonals distributed in interval $[-r,r]$ with a common difference $a\geq 1$, and an algorithm $f$, we say $f$ is an ideal $-\ell$ decomposition ($1\leq\ell<\log{(r/a)}$) of $U$ if $f(U,\ell)=U_{L_{\ell}}U_{R_{\ell}}U_{R_{\ell-1}}\dots U_{R_{1}}$ satisfies:
  1. (1)
    For any $U_{R_{i}},1\leq i\leq\ell$, it has non-zero diagonals only at diagonals with indices $\{0,\pm\lceil r_{c}^{(i)}\rceil\}$ where $r_{c}^{(i)}=a\cdot\lceil\frac{r/a}{2^{i}}\rceil$.
  2. (2)
    The non-zero diagonals of $U_{L_{\ell}}$ are distributed between $[-r^{\prime},r^{\prime}]$, where $r^{\prime}=r-\sum_{i=1}^{\ell}r_{c}^{(i)}$.
If $f$ is an ideal depth-$\log{(r/a)}-1$ decomposition, then we may also call it an ideal full-depth decomposition.
However, intuitively, not all permutations possess such an ideal decomposition, and it is even non-trivial to check whether a permutation has one. To our knowledge, the most relevant work is that of Han et al. (Han et al., 2019), which only proposes a decomposition of the discrete Fourier transformation (DFT) satisfying Definition 2 rather than any pure permutation.
3. Searching Ideal Depth-1 Decomposition
To find a full-depth ideal (or approximate ideal) decomposition for a given permutation, we can first search for a depth-1 solution and see if some pattern exists for constructing a $-\ell$ solution for $\ell>1$. This section aims to complete the first step, designing an algorithm to search for a depth-$1$ ideal decomposition of any permutation.
3.1. Basic Decomposition Pattern
To factorize an $n\times n$ permutation matrix $U$ into a product of two factors $U=U_{L}U_{R}$ satisfying the ideal depth-$1$ decomposition of Definition 2, we first observe how non-zero entries on diagonals of $U_{L}$ and $U_{R}$ contribute to those in $U$. For any non-zero entry $U(k,l)_{r}$ in $U$, it is determined by the $k_{1}$-th diagonal in $U_{R}$ and the $k_{2}$-th diagonal in $U_{L}$, where $k_{2}=k-k_{1}\bmod n$ (Theorem 1). This follows that $U_{R}(k_{1},k+l-k_{1})_{r}$ and $U_{L}(k_{2},l)_{r}$ should be set to $1$ to guarantee the correct generation of $U(k,l)_{r}$. We call this decomposing $U(k,l)$ into $k_{1}$-th diagonal and $k_{2}$-th diagonal (see Figure 1). Though entry $U(k,l)$ can be associated with more than two diagonals in general, it will end up with $U_{R}$ and $U_{L}$ not being permutations, which may damage the composability.
In the following context, we use function ConvertR to describe the operation of decomposing any entry $U(k,l)_{r}$ into the $k_{1}$-th diagonal of $U_{R}$:
(4) $\texttt{ConvertR}(k,l,k_{1})=k+l-k_{1}.$
The output is the row index of the entry in $U_{R}$ that will be set to $1$.
Figure 1. Decompose single entry in $8\times 8$ permutation matrix
[…]
[Decomposing single entry in an $8\times 8$ permutation matrix]
During decomposing entries of $U$ into some desired diagonals of $U_{R}$ and $U_{L}$, conflicts (Definition 1) may occur and damage the correctness of factorization. In $U_{R}$, row conflicts may occur, while column conflicts are impossible; conversely, the opposite holds in $U_{L}$. Additionally, since the column index of the entry to be set to $1$ in $U_{L}$ equals $\texttt{ConvertR}(k,l,k_{1})$, a duality exists between the two matrices, indicating that managing either $U_{L}$ or $U_{R}$ is sufficient to make the other matrix undergo the same process.
Definition 0.
We say there is a row conflict (or column conflict) between two non-zero entries in a permutation matrix, if they are in the same row (or column), and there is a conflict between two non-zero diagonals if one or more pairs of conflicting entries exist on these diagonals.
Algorithm 1 Check One Conflicting Row Between the 00-th and ±$\pm r_{c}$-th Diagonals (CheckRow)
1:
2:$M=\{M_{r_{c}},M_{-r_{c}},M_{0}\}$: Non-zero diagonals in $U_{R}$;
3:$n$: The dimension of $U$;
4:$r_{c}$: The non-zero diagonal index of $U_{R}$;
5:$a_{0}$: The row with a potential conflict to be resolved;
6:$Q$: The set containing indices of rows where conflicts exist;
7:$i$: The number of resolved rows in $Q$.
8:
9:$solved\in\{0,1\}$
10:if $M_{0}[a]=\varnothing$ then $\triangleright$ Row $a$ has no conflict
11:     if $i+1<\textsc{len}(Q)$ then $\triangleright$ Find next conflict to resolve
12:         $i\leftarrow i+1,a\leftarrow Q[i]$
13:     $\triangleright$ All conflicts resolved
14:         return $1$
15:     end if
16:end if
17:[Try to move $U_{R}(0,a)_{r}$ to the $r_{c}$-th diagonal]
18:$b\leftarrow M_{0}[a],M_{0}[a]\leftarrow\varnothing,\hat{a}\leftarrow(a-r_{c})\bmod{n},solved\leftarrow 0$
19:if $b\geq 0$ and $M_{-r_{c}}[\hat{a}]=\varnothing$ then
20:     $M_{r_{c}}[\hat{a}]\leftarrow b$
21:     $solved\leftarrow\texttt{CheckRow}(M,n,r_{c},\hat{a},Q,i)$triangleright Resolve new conflicts caused by the movement
22:end if
23:[Otherwise, try to move $U_{R}(0,a)_{r}$ to the $-r_{c}$-th diagonal]
24:if $solved=0$ and $b\leq 0$ then
25:     $M_{r_{c}}[\hat{a}]\leftarrow\varnothing,\hat{a}\leftarrow a+r_{c}\bmod{n}$
26:     if $M_{r_{c}}[\hat{a}]=\varnothing$ then
27:         $M_{-r_{c}}[\hat{a}]\leftarrow b$
28:         $solved\leftarrow\texttt{CheckRow}(M,n,r_{c},\hat{a},Q,i)$triangleright Resolve new conflicts caused by the movement
29:     end if
30:end if
31:if $solved=0$ then $\triangleright$ Cannot resolve, undo modifications
32:     $M_{r_{c}}[\hat{a}]\leftarrow\varnothing,M_{-r_{c}}[\hat{a}]\leftarrow\varnothing,M_{0}[a]\leftarrow b$
33:end if
34:return $solved$triangleright Return to the previous recursive call
3.2. Concrete Design
Based on the observation above, we design an algorithm searching a depth-1 ideal decomposition $U=U_{L}U_{R}$ for arbitrary $n\times n$ permutation $U$. The algorithm is composed of three steps.
Step 1. (Decompose Entries) This step decomposes each non-zero entry in $U^{\prime}$ to the $r_{c}$-th, $-r_{c}$-th, or 00-th diagonals of $U_{R}$ where $r_{c}$ corresponds to $r_{c}^{(1)}$ in Definition 2). First, we create a set of key-value tables representing the non-zero diagonals of $U_{R}$ :
$M=\{M_{r_{c}},M_{-r_{c}},M_{0}\},$
where $M_{i}$ represents the $i$-th non-zero diagonal for $i\in\{\pm r_{c},0\}$. For each table, its keys indicate the row coordinates of entries on it, and its values denote the diagonal coordinates of entries in $U$ determined by the keys. Next, for each non-zero entry $U(k,l)_{r}$ in $U^{\prime}$:
  1. (1)
    If $k\geq r_{c}$ and $M_{-r_{c}}[\texttt{ConvertR}(k,l,r_{c})]$ is empty, add the key-value pair $\langle\texttt{ConvertR}(k,l,r_{c}),k\rangle$ to $M_{r_{c}}$.
  2. (2)
    If $k\leq-r_{c}$ and $M_{r_{c}}[\texttt{ConvertR}(k,l,-r_{c})]$ is empty, add the key-value pair $\langle\texttt{ConvertR}(k,l,-r_{c}),k\rangle$ to $M_{-r_{c}}$.
  3. (3)
    If $-r_{c}<k<r_{c}$, add $\langle\texttt{ConvertR}(k,l,0),k\rangle$ to $M_{0}$;
  4. (4)
    Otherwise, a conflict that cannot be resolved exists, and the algorithm terminates.
The explanation of the conditional statements above is given as follows:
According to Theorem 1 and Definition 2, for the diagonal distribution of $U_{L}$ to converge to $[-r+r_{c},r-r_{c}]$ with $r_{c}=a\cdot\lceil\frac{r/a}{2}\rceil$, non-zero entries on the $k$-th diagonal of $U$ with $|k|>r_{c}$ can only be decomposed to the $\texttt{sign}(k)\cdot r_{c}$-th diagonal of $U_{R}$. Step 1 employs the first two conditional statements to receive these entries for decomposition. If conflicts arise between the $r_{c}$-th and $-r_{c}$-th diagonals during the decomposition of these entries, it indicates that no ideal decomposition exists for $U$. Such cases are captured by the fourth conditional statement.
The remaining non-zero entries of $U$ are distributed on diagonals whose indices’ absolute value do not exceed $r_{c}$, and these entries can either be decomposed to the 0-th diagonal or the nearest diagonal among the $r_{c}$-th and $-r_{c}$-th diagonals according to Theorem 1. The strategy in Step 1 is to decompose all of them to the 0-th diagonal.
After Step 1, $U_{R}$ has non-zero entries only on the 0-th and ±$\pm r_{c}$-th diagonals, with potential conflicts occurring only between the 0-th and $r_{c}$-th diagonals, as well as between the 0-th and $-r_{c}$-th diagonals.
Step 2. (Resolve Conflicts) This step aims to resolve the conflicts derived from the previous step. We first create a set $Q$ to record the rows where conflicts exist. For each key-value pair $\langle a_{1},b_{1}\rangle$ in $M_{0}$, if there exists another key-value pair $\langle a_{1},b_{2}\rangle$ in $M_{r_{c}}$ or $M_{-r_{c}}$, it indicates a conflict in row $a_{1}$ of $U_{R}$, and we add $a_{1}$ to the set $Q$. Then, we compute $solved\leftarrow\texttt{CheckRow}(M,n,r_{c},Q[0],Q,0)$ (Algorithm 1). If $solved=1$, all the conflicts are successfully resolved; otherwise, there are irresolvable conflicts, and the algorithm terminates.
Assigned with some row index $a$, the function CheckRow first checks if any conflict exists in row $a$. If there is a conflict, it attempts to resolve the conflict; otherwise, it proceeds to resolve the next conflict in $Q$. From the previous step, it is known that the conflicts occur between the 0-th diagonal and the ±$\pm r_{c}$-th diagonals. Therefore, for any conflicting row $a$, CheckRow can only resolve it by moving the entry $U_{R}(0,a)_{r}$ to the $r_{c}$-th (or $-r_{c}$-th) diagonal. The condition for successfully moving $U_{R}(0,a)_{r}$ is that the target row $\hat{a}$ must not contain any entries on the $r_{c}$-th (or $-r_{c}$-th) diagonal. If such an entry exists (with a corresponding key-value pair $\langle\hat{a},v\rangle$ in $M_{k}$ for $k=r_{c}$ or $-r_{c}$), two possible cases arise:
  1. (1)
    If $|v|>r_{c}$, then $U_{R}(0,a)_{r}$ conflicts with $U_{R}(k,\hat{a})_{r}$ after the movement, but $U_{R}(k,\hat{a})_{r}$ is not movable.
  2. (2)
    Otherwise, $U_{R}(k,\hat{a})_{r}$ must have originally been moved from the 00-th diagonal to row $\hat{a}$. If continuing to move this entry results in an ideal decomposition, by Theorem 2, there must be an alternative order of entry movements that avoids moving any entry more than once and still reaches the same solution.
Therefore, $U_{R}(0,a)_{r}$ should not be moved to row $\hat{a}$ in either of these cases. If the movement is successful, CheckRow invokes itself to resolve any potential conflict between the $r_{c}$-th (or $-r_{c}$-th) diagonal and the 00-th diagonal for row $\hat{a}$. If the invocation returns 00 for irresolvable conflicts, then CheckRow has to undo the movement it employs and return 00 to the previous recursive call. If the invocation determines no conflict exists in row $\hat{a}$, it proceeds to resolve the next conflict recorded in $Q$. In this way, CheckRow operates as a depth-first search, exhaustively checking all entry movement combinations until it identifies the first conflict-free $U_{R}$.
Theorem 2.
For any depth-1 ideal decomposition $U=U_{L}^{\prime}U_{R}^{\prime}$ of $U$, there exists an order of entry movements where each entry in $U_{R}$ is moved at most once to obtain $U_{R}^{\prime}$ after Step 1 of the search algorithm.
Proof.
For $0\leq i<n$, move the non-zero entry in the $i$-th column of $U_{R}$ to the position of the non-zero entry in the $i$-th column of $U_{R}^{\prime}$. Clearly, each entry is moved only once. ∎
Step 3. (Reconstruct $U_{L}$). Generate $U_{L}$ according to $U$ and $U_{R}$ such that $U=U_{L}U_{R}$. One may replace $U$ with $U_{L}$ and update $r_{c}$ for depth-$2$ ideal decomposition search.
The correctness of the search algorithm that it detects whether the given permutation $U$ has any ideal depth-1 decomposition is demonstrated throughout the description above. In the next section, the search algorithm is applied to specific permutations in homomorphic matrix operations, where we provide a visualizable demonstration of this algorithm.
4. Full-depth Ideal Decomposability on Specific Permutations
In this section, we demonstrate that certain permutations can be ideally decomposed in full depth as per Definition 2. We first apply the depth-$1$ decomposition searching algorithm from Section 3 to the permutation used in Jiang et al.’s HMT (Jiang et al., 2018), through which we discover the full-depth ideal decomposability of this permutation. We further reduce the permutations from Rizomiliotis and Triakosia’s HMM (Rizomiliotis and Triakosia, 2022) to the former, illustrating their decomposability.
4.1. Homomorphic Matrix Transposition
Let $ct(\mathbf{a})$ denote a ciphertext of an $n$-dimensional vector $\mathbf{a}$, where $\mathbf{a}$ is the row-ordering encoding of a $d\times d$ matrix $A$ ($d^{2}\leq n$). To obtain the ciphertext of the row-ordering vector of $A^{T}$, a specific homomorphic permutation (denoted as $n\times n$ matrix $U^{t}$) is applied to $ct(\mathbf{a})$. $U^{t}$ possesses $12d-$ 1 nonzero diagonals defined as:
$U^{t}\Big{(}i(d-1),j(d+1)+ik\Big{)}_{r}=\begin{cases}1,&i\in[0,d),k=1;\\ 1,&i\in(-d,0),k=-d;\\ 0,&\text{otherwise}.\\ \end{cases}$
where $j\in[0,d-|i|)$. Currently, the BSGS-based HLT mentioned previously yields the best performance for $U^{t}$ with an overhead of approximately $d_{1}+2d_{2}$ rotations and keys, where $d_{1}d_{2}=d$. As noted previously, using the Benes network for decomposition would yield a rotation complexity no less than $O(4\log{d^{2}})$, which is inefficient within the range $0<d<2^{8}$ for practical HE parameter settings (Bossuat et al., 2024), not to mention the required $2\log{d^{2}}$ multiplicative depth.
4.1.1. Case with Dimension a Power of $2$
For $d$ a power of $2$, $U^{t}$ can be precisely decomposed to $U^{t}_{L}U^{t}_{R}$ when we apply the depth-$1$ decomposition searching algorithm proposed in Section 3. Figure 2 illustrates this process of a $16\times 16$ $U^{t}$.
The $r_{c}$ is set to $(d-1)\cdot\lceil\frac{d-1}{2}\rceil=6$ as the indices of non-zero diagonals in $U^{t}$ have a common difference of $d-1=3$. Step 1 of the algorithm is presented in Figure 2b. We can observe that $U^{t}$’s entries in diagonals $k\geq 6$ and $k\leq-6$ are decomposed to the $6$-th and $-6$-th diagonal of $U_{R}^{t}$ respectively, while those in diagonal $-6<k<6$ are decomposed to the 00-th diagonal. This introduces conflicts in rows $6$ and $9$. Figure 2c shows how Step 2 resolves the conflicts, illustrated by the dash paths. For example, let CheckRow begin with row $9$: CheckRow first moves the conflicting entry $U_{R}(0,9)$ to the $6$-th diagonal as $M_{0}[9]>0$, and this corresponds to a destination of row $3$. CheckRow then invokes itself to check row $3$. Determining that no conflict exists at row $3$, the invocation proceeds to resolve the conflict at row $6$.
Once we solve all the conflicts in $U^{t}_{R}$, we recover $U^{t}_{L}$ such that $U^{t}=U^{t}_{L}U^{t}_{R}$ (Figure 2d). We can observe that both of these factors exhibit good structural properties.
(a) $16\times 16$ $U^{t}$ with non-zero entries painted in dark blue.
(b) Step 1. Red entries denote the non-zero entries in $U^{t}_{R}$ after Step 1.
(c) Step 2. $U_{R}^{t}$’s final view. Conflicting entries moved away are painted in grey
(d) Step 3. $U^{t}_{L}$ reconstructed by $U^{\sigma}_{R}$ and $U^{t}$.
Figure 2. Applying depth-1 ideal decomposition search to $U^{t}$ of size $16\times 16$.
If we let $A$ be evenly partitioned into $4$ blocks with size $d/2\times d/2$:
(5) $A=\left[\begin{matrix}{A^{\prime}_{0,0}}&{A^{\prime}_{0,1}}\\ {A^{\prime}_{1,0}}&{A^{\prime}_{1,1}}\\ \end{matrix}\right],$
then the effect of $U^{t}_{R}$ and $U^{t}_{L}$ can be expressed as follows:
(6) $\left[\begin{matrix}{A^{\prime}_{0,0}}&{A^{\prime}_{1,0}}\\ {A^{\prime}_{0,1}}&{A^{\prime}_{1,1}}\\ \end{matrix}\right]\leftarrow U^{t}_{R}(A),\left[\begin{matrix}{A^{\prime}_{0,0}}^{T}&{A^{\prime}_{1,0}}^{T}\\ {A^{\prime}_{0,1}}^{T}&{A^{\prime}_{1,1}}^{T}\\ \end{matrix}\right]\leftarrow U^{t}_{L}\left(U^{t}_{R}(A)\right).$
Equation 6 indicates that $U^{t}_{R}$ swaps the positions of $A^{\prime}_{0,1}$ and $A^{\prime}_{1,0}$, while $U^{t}_{L}$ performs block-wise transposition. This follows that:
(7) $A^{T}=U_{L_{\ell}}^{t}\left(U_{R_{\ell}}^{t}\left(\dots\left(U_{R_{1}}^{t}(A)\right)\right)\right)$
holds for $\ell\leq\log{d}-1$ if the following process defines the permutations in it:
  1. (1)
    $U^{t}_{R_{i}}$ evenly partitions each block of $A$ into $4$ sub-blocks (with $A$ itself being the initial block). The sub-blocks are interpreted as blocks in the next partition before $U^{t}_{R_{i+1}}$.
  2. (2)
    $U^{t}_{R_{i}}$ swaps the sub-blocks at positions $[0,1]$ and $[1,0]$ in each block of $A$.
  3. (3)
    $U^{t}_{L_{\ell}}$ performs block-wise transposition.
These rules indicate that $U^{t}_{R_{i}}$ swaps $d/2^{i}$-size sub-blocks inside each $d/2^{i-1}$-size block, where its non-zero diagonals are precisely distributed at $\{0,\pm(d-1)d/2^{i}\}$ according to Theorem 1 and Theorem 2. On the other hand, $U^{t}_{L_{\ell}}$ transposes each $d/2^{\ell}$-size block by swapping their entries at positions $[q,p]$ and $[p,q]$ for $0\leq p<q<d/2^{\ell}$. These swappings correspond to non-zero diagonals with indices:
{±$\begin{split}&\{\pm(d(q-p)+p-q)\mid 0\leq p<q<d/2^{\ell})\}=\\ &\{\pm i\cdot(d-1)\mid 0\leq i<d/2^{\ell}\}.\end{split}$
Now, it is evident that Equation 7 with $U^{t}_{R_{i}}$ and $U^{t}_{\ell}$ defined above is a $-\ell$ ideal decomposition of $U^{t}$ for $11\leq\ell\leq\log{d}-$ 1.
Theorem 1.
The permutation matrix swapping any two blocks of any matrix contains at most three nonzero diagonals.
Proof.
Given $A$ is of size $c\times d$, and the two same size blocks to be swapped are $A^{\prime}_{i,j}$ and $A^{\prime}_{k,l}$. If $A^{\prime}_{i,j}[0,0]=A[p,q]$ and $A^{\prime}_{k,l}[0,0]=A[u,v]$, for any two entries $A^{\prime}_{i,j}[x,y]$ and $A^{\prime}_{k,l}[x,y]$, we have:
(8) $\begin{split}A^{\prime}_{i,j}[x,y]=A[p+x,q+y],\\ A^{\prime}_{k,l}[x,y]=A[u+x,v+y].\end{split}$
Let $r=(u+x)d+(v+y)-(p+x)d-(q+y)=d(u-p)+(v-q)$. Moving $A^{\prime}_{i,j}$ to the position of $A^{\prime}_{k,l}$ requires a rotation by step $-r$, while moving $A^{\prime}_{k,l}$ to the position of $A^{\prime}_{i,j}$ requires rotation by step $r$. Entries outside the blocks maintain their original positions requiring rotation by step 00. Furthermore, if either $A^{\prime}_{i,j}$ or $A^{\prime}_{k,l}$ is entirely zero, the corresponding rotation for that block can be omitted. Thus, the swapping involves at most three non-zero diagonals in its matrix representation. ∎
Theorem 2.
The permutations that perform the same operation on different blocks of the same size within a matrix have identical non-zero diagonal distributions.
Proof.
Given $A$ is of size $c\times d$, let $A^{\prime}$ be any submatrix of $A$ with $A^{\prime}[0,0]=A[s,t]$. For any two entries $A^{\prime}[p,q]$ and $A^{\prime}[u,v]$, we have:
(9) $\begin{split}A^{\prime}[p,q]=A[s+p,t+q],\\ A^{\prime}[u,v]=A[s+u,t+v].\end{split}$
To move $A^{\prime}[p,q]$ to the position of $A^{\prime}[u,v]$, it requires a rotation of $d(s+u)+(t+v)-d(s+p)-(t+q)=d(p-u)+(q-v)$ steps. It is evident that the number of different rotations for movements within the block is independent of the block’s position within $A$. ∎
4.1.2. Case with Arbitrary Dimension
We observe that when the search algorithm is recursively applied to $U^{t}$ with arbitrary dimension $d$, it always terminates after $k$ recursion, where $k$ is the largest power of $2$ dividing $d$. This aligns with the decomposition rule for power-of-2 cases we proved above: $U^{t}_{R_{k+1}}$ is not defined as the blocks in $A$ cannot be evenly partitioned into $4$ sub-blocks at that point.
The simplest solution is to pad $A$ with zeros, increasing its dimensions to the nearest power of 2, $2^{i}$. Then the corresponding $U^{t}$ can undergo an ideal decomposition of depth $\lceil\log{d}\rceil-1$. This approach is practical unless $(2^{i})^{2}$ exceeds the maximum encoding length $n$ of the ciphertext. We propose an alternative method to avoid this drawback. The idea is to define a new method for $U^{t}_{R_{i}}$ to partition and swap odd-dimensional blocks. Specifically, for any $d\times d$ block $A$ with $d=2\cdot d^{\prime}+1$ for some $0\leq d^{\prime}$, let:
(10) $\begin{array}[]{ll}B^{\prime}_{0,0}=\begin{bmatrix}A[i,j]\end{bmatrix}_{0\leq i\leq d^{\prime}}^{0\leq j\leq d^{\prime}}&B^{\prime}_{0,1}=\begin{bmatrix}A[i,j]\end{bmatrix}_{0\leq i\leq d^{\prime}-1}^{d^{\prime}+1\leq j\leq 2d^{\prime}}\\ \\ B^{\prime}_{1,0}=\begin{bmatrix}A[i,j]\end{bmatrix}_{d^{\prime}+1\leq i\leq 2d^{\prime}}^{0\leq j\leq d^{\prime}-1}&B^{\prime}_{1,1}=\begin{bmatrix}A[i,j]\end{bmatrix}_{d^{\prime}\leq i\leq 2d^{\prime}}^{d^{\prime}\leq j\leq 2d^{\prime}}\end{array}.$
The partition of $A$ into unequal blocks is given as:
(11) $A=\begin{bmatrix}B^{\prime}_{0,0}&B^{\prime}_{0,1}\\ B^{\prime}_{1,0}&B^{\prime}_{1,1}\end{bmatrix},$
with an overlapping part $A[d^{\prime},d^{\prime}]$ between $B^{\prime}_{0,0}$ and $B^{\prime}_{1,1}$. In any partitioning round, the blocks have at most two dimensions differing by at most 1 (proof given in Appendix B.1). In this case, the non-zero diagonals in the block-wise permutation are the union of those required for the permutations on blocks of different sizes. Thus, for $1<i\leq\ell$, the maximum number of non-zero diagonals in each $U^{t}_{R_{i}}$ increases from 3 to 5. The number of non-zero diagonals of $U^{t}_{L_{\ell}}$ remains $O(d/2^{\ell-1})$ for $\ell\leq\lceil\log{d}\rceil-1$. Such a decomposition closely approximates the ideal form and avoids the possible increment of $n$ from the zero-padding strategy.
4.2. Homomorphic Matrix Multiplication
It is intriguing that the permutations in Rizomiliotis and Triakosia’s HMM algorithm (Rizomiliotis and Triakosia, 2022) have decomposability similar to $U^{t}$. We can achieve an asymptotically faster HMM by applying decompositions to these permutations. Furthermore, we present methods to reduce the multiplication depth incurred in the decomposition to a constant level and to enhance the utilization of the encoding space.
Let $A$ and $B$ be the $d\times d$ matrices to be multiplied, and let the plaintext space be a vector space of dimension $n=d^{3}$. The method proposed by Rizomiliotis and Triakosia first appends $(d-1)$ zero matrices of size $d\times d$ to both the end of $A$ and $B$ to form $\hat{A}$ and $\hat{B}$ of size $d^{2}\times d$, where each of them can be encrypted by a ciphertext. Two permutations $U^{\gamma}$ and $U^{\xi}$ are then applied to $\hat{A}$ and $\hat{B}$ respectively. Next, column-wise replication is performed on $U^{\gamma}(\hat{A})$ and row-wise replication on $U^{\xi}(\hat{B})$. Denote the replicated results as $\tilde{A}$ and $\tilde{B}$. Finally, compute $\hat{C}=\tilde{A}\odot\tilde{B}$, and perform $\hat{C}\leftarrow\texttt{Rot}(\hat{C},d^{3}/2^{i})+\hat{C}$ for $1\leq i\leq\log{d}$. This yields the desired result $C=AB$ in the first $d\times d$ block of $\hat{C}$.
Compared to previous HMM algorithms (Halevi and Shoup, 2018; Jiang et al., 2018), Rizomiliotis and Triakosia’s approach is characterized by its novel use of a $d^{3}$ redundant encoding space, which minimizes the number of ciphertext multiplications and rotations. Its complexity is considered to be dominated by the $U^{\gamma}$ and $U^{\xi}$ permutations, which necessitate $2d$ ciphertext rotations.
4.2.1. Decomposability of $U^{\gamma}$ and $U^{\xi}$
We observe that $U^{\gamma}$ can be fully regarded as a transposition acting on a $d\times d$ matrix where each entry is a $d\times 1$ vector:
(12) $\begin{split}&\hat{A}=\left[\begin{matrix}\mathbf{a}_{0}&\cdots&\mathbf{a}_{d-1}\\ \mathbf{0}_{c}&\cdots&\mathbf{0}_{c}\\ \vdots&\ddots&\vdots\\ \mathbf{0}_{c}&\cdots&\mathbf{0}_{c}\\ \end{matrix}\right],U^{\gamma}(\hat{A})=\left[\begin{matrix}\mathbf{a}_{0}&\mathbf{0}_{c}&\cdots&\mathbf{0}_{c}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{a}_{d-1}&\mathbf{0}_{c}&\cdots&\mathbf{0}_{c}\\ \end{matrix}\right],\\ &\mathbf{a}_{i}=\left[\begin{matrix}A[0,i]&\cdots&A[d-1,i]\end{matrix}\right]^{T},\text{for}\phantom{a}0\leq i<d.\end{split}$
The rule from the previous section can be applied to this transposition. Thus, the decomposition of $U^{\gamma}$ is similar to $U^{t}$. Specifically, when $d$ is a power of two, $U^{\gamma}$ has a $-\ell$ ideal decomposition for $\ell\leq\log{d}-1$. When $d$ is set arbitrarily, $U^{\gamma}$ has an approximate $-\ell$ ideal decomposition for $\ell\leq\lfloor\log{d}\rceil-1$.
Furthermore, we can modify the decomposition of $U^{\gamma}$ to achieve constant multiplication depth. We fill the non-zero diagonals of $U^{\gamma}$ with $1$ and denote the result as $U^{\gamma}_{pad}$ (proof provided in Appendix B.2). At this point, $U^{\gamma}_{pad}(\hat{A})$ is the sum of the rotated $\hat{A}$ by steps of $0,-(d^{2}-1),\dots,-(d^{2}-1)(d-1)$ without any masking. It can be observed that $U^{\gamma}_{pad}(\hat{A})$ has the same first column as $U^{\gamma}(\hat{A})$. This implies that we can get $U^{\gamma}(\hat{A})$ by masking $U^{\gamma}_{pad}(\hat{A})$ only once. The decomposition of $U^{\gamma}_{pad}$ can be directly obtained by filling all non-zero diagonals in the permutation matrices of $U^{\gamma}$’s decomposition with $1$, yielding a constant multiplication depth of $1$.
Similarly, $U^{\xi}$ can be interpreted as a transposition acting on a $d\times d$ matrix where each basic unit is a $1\times d$ vector:
(13) $\begin{split}&\hat{B}=\left[\begin{matrix}\mathbf{b}_{0}&\cdots&\mathbf{b}_{d-1}\\ \mathbf{0}_{r}&\cdots&\mathbf{0}_{r}\\ \vdots&\ddots&\vdots\\ \mathbf{0}_{r}&\cdots&\mathbf{0}_{r}\\ \end{matrix}\right],U^{\xi}(\hat{B})=\left[\begin{matrix}\mathbf{b}_{0}&\mathbf{0}_{r}&\cdots&\mathbf{0}_{r}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{b}_{d-1}&\mathbf{0}_{r}&\cdots&\mathbf{0}_{r}\\ \end{matrix}\right],\\ &\mathbf{b}_{i}=\left[\begin{matrix}B[i,0]&\cdots&B[i,d-1]\end{matrix}\right],\text{for}\phantom{a}0\leq i<d.\end{split}$
Thus, $U^{\xi}$ can be decomposed following the decomposition method of $U^{t}$, and its non-zero diagonals can also be padded with $1$ to achieve a constant multiplication depth.
4.2.2. Flexible Utilization of the Encoding Space
The original algorithm by Rizomiliotis and Triakosia necessitates a plaintext encoding space of $d^{3}$ to perform a $d\times d$ matrix multiplication, resulting in a fast-growing computational overhead as $d$ increases. Here, we propose a strategy to enable configurable encoding space. In essence, we split $U^{\gamma}(\hat{A})$ and $U^{\xi}(\hat{B})$ each into $d/d^{\prime}$ blocks (each size is $d^{\prime}\times d^{\prime}$ vectors), reducing the encoding space required for representing $A$ and $B$ to $d^{2}\cdot d^{\prime}$. Let $d^{\prime}$ be a divisor of $d$, $\hat{A}$ and $\hat{B}$ are both interpreted as the concatenation of $d/d^{\prime}$ blocks as follows:
(14) $\hat{G}$ $=\left[\begin{array}[]{c|ccc|c}\cdots&\mathbf{g}_{d^{\prime}k}&\cdots&\mathbf{g}_{d^{\prime}(k+1)-1}&\cdots\\ \cdots&\mathbf{0}&\cdots&\mathbf{0}&\cdots\\ \cdots&\vdots&\ddots&\vdots&\cdots\\ \cdots&\mathbf{0}&\cdots&\mathbf{0}&\cdots\\ \end{array}\right]$
$=\left[\begin{matrix}\hat{G}^{\prime}_{0}&\hat{G}^{\prime}_{1}&\cdots&\hat{G}^{\prime}_{\frac{d}{d^{\prime}}-1}\end{matrix}\right],{(0\leq k<\frac{d}{d^{\prime}})}$
where $\hat{G}=\hat{A},\mathbf{g}_{i}=\mathbf{a}_{i},\mathbf{0}=\mathbf{0}_{c}$ (or $\hat{G}=\hat{B},\mathbf{g}_{i}=\mathbf{b}_{i},\mathbf{0}=\mathbf{0}_{r}$). To complete an HMM, we apply block-wise $U^{\gamma}$ and $U^{\xi}$ to each $\hat{A}^{\prime}_{i}$ and $\hat{B}^{\prime}_{i}$ respectively, yielding: $U^{\gamma^{\prime}}(\hat{A})=[U^{\gamma}(\hat{A}^{\prime}_{0})\ldots U^{\gamma}(\hat{A}^{\prime}_{d/d^{\prime}-1})]$ and $U^{\xi^{\prime}}(\hat{B})=[U^{\xi}(\hat{B}^{\prime}_{0})\ldots U^{\xi}(\hat{B}^{\prime}_{d/d^{\prime}-1})]$. This costs $2\log{d^{\prime}}$ rotations in total (Theorem 2). For each $U^{\gamma}(\hat{A}^{\prime}_{i})$ (or $U^{\xi}(\hat{B}^{\prime}_{i})$), we use masking to extract it then replicate its first colum to the entire encoding space individually, yielding the output $\tilde{A}_{i}$ (or $\tilde{B}_{i}$). This costs $2\frac{d}{d^{\prime}}\log{d}$ rotations in total. Finally, we compute $\hat{C}=\sum_{i=0}^{\frac{d}{d^{\prime}}-1}\tilde{A}_{i}\odot\tilde{B}_{i}$, followed by a summation over $\log d^{\prime}$ rotations on $\hat{C}$ to output the desired result. The total number of rotations is $3\log d^{\prime}+2\frac{d}{d^{\prime}}\log d$.
On the other hand, a ciphertext may store more matrices as the encoding space requirement decreases. For $m$ pairs of $d\times d$ matrices $A_{i}$ and $B_{i}$ ($0\leq i<m$) stored in $\hat{\mathcal{A}}=[\hat{A}_{0},\ldots,\hat{A}_{m-1}]^{T}$, the HMM between $\hat{\mathcal{A}}$ and $\hat{\mathcal{B}}$ yields $\hat{\mathcal{C}}=[\hat{C}_{0},\ldots,\hat{C}_{m-1}]^{T}$, where each $\hat{C}_{i}$ contains $C_{i}=A_{i}B_{i}$. Such a parallelizable setup gives an amortized rotation complexity of $\left(\frac{3}{m}\log d^{\prime}+2\frac{d}{d^{\prime}m}\log d\right)$. Further trade-offs are discussed in Appendix B.3 for reducing the complexity component $(\frac{d}{d^{\prime}}\log{d})$ of replications to $\left(\frac{d}{d_{0}}+\frac{d}{d^{\prime}}\log{d_{0}}\right)$ for $1\leq d_{0}\leq d$.
Scheme Rot Mult Depth Space Parallelity
(Jiang et al., 2018) $3d+5\sqrt{d}$ $d$ $3$ $d^{2}$ $\frac{n}{d^{2}}$
(Rizomiliotis and Triakosia, 2022) $2d+3\log d$ $1$ $2$ $d^{3}$
(Huang and Zong, 2023) $d\log{d}+d$ $d$ $2$ $d^{2}$
(Zhu et al., 2023) $4d$ $2d$ $2$ $d^{2}$
Ours $2d\left(\frac{1}{d_{0}}+\frac{\log{d_{0}}}{d^{\prime}}\right)+3\log{d^{\prime}}$ $\frac{d}{d^{\prime}}$ 2 $d^{2}d^{\prime}$ $\frac{n}{d^{2}d^{\prime}}$
  • Parallelity is not explicitly discussed in some schemes.
Table 1. Complexity of HMMs under row/column-major encoding
4.2.3. Complexity Comparison
Table 1 presents existing HMMs under row (or column)-major encoding, where matrices are mapped to the plaintext (vector) space of the batch-encoding HE schemes in a row-wise or column-wise unfolded manner. Jiang et al.’s scheme (Jiang et al., 2018) achieves optimal complexity when considering amortized runtime. Rizomiliotis and Triakosia (Rizomiliotis and Triakosia, 2022) achieve a reduced number of Mults and Rots with a $d$-fold encoding space. However, the increased encoding space may lead to expanded ciphertexts with higher latency per basic operation. Huang et al. (Huang and Zong, 2023) and Zhu et al. (Zhu et al., 2023) design HMMs for rectangular matrices but with modest performance for square matrices. They also require a hypercube packing method supported by BGV-typed HE schemes(Gentry et al., 2012; Brakerski et al., 2014), lacking compatibility with the one-dimensional encoding space. Our method does not require a hypercube structure, so it is compatible to all batch-encoding HE schemes. We also offer configurability in terms of complexity, encoding space, and rotation keys. Notably, we only needs $5\log{n}$ Rots for the maximum redundant encoding space: $(d^{\prime}=d,d_{0}=d)$, and $2d$ Rots for the minimum encoding space: $(d^{\prime}=1,d_{0}=1)$.
At the end of this section, it is worth noting that the HMT and HMM optimized above do not represent the entirety of the depth-1 ideal decomposition search algorithm. Due to space limitations, more applications of the search algorithm are presented in Appendix C, where we use it to find the full-depth decomposability of another two important permutations used in Jiang et al.’s HMM (Jiang et al., 2018).
5. New Design for Arbitrary Permutation Decomposition
The idea of ideal decomposition aims to establish an upper bound for decomposition techniques. As previously discussed, this bound can indeed be achieved for specific permutations. However, not all permutations can be ideally decomposed. For example, the search algorithm struggles to find solutions for randomly generated permutations.
The Benes network-based decomposition (Gentry et al., 2012) combined with its optimal level-collapsing strategy proposed by Halevi and Shoup (Halevi and Shoup, 2014) may form a decomposition chain with multiplication depth $\log{n}-1$ for any given permutation with length $n$, where each factor contains approximately 6 to 7 non-zero diagonals on average. Although this already offers an approximation to ideal decomposition, we aim to present a new solution outperforming such an approximation in certain scenarios.
Figure 3. Proposed multi-group network design for arbitrary homomorphic permutation (using permutation of length 16 as an example)
5.1. Overall Design
Our design conceptualizes permutations as a directed graph network (Figure 3). It consists of four major components:
  • Nodes. The nodes in the network are categorized into two types: rotation nodes and standby nodes, both taking vectors as input. Rotation nodes rotate the input vector by certain steps, while standby nodes output the input vector unchanged.
  • Edges. An edge applies a mask to the output of a node and transmits it to another node. When multiple edges converge at the same node, the vectors they carry are summed together to form the input to that node.
  • Levels. Levels correspond to the sequence in which nodes are processed. Nodes at the same level are handled concurrently.
  • Groups. The network is vertically divided into several groups, each containing a column of rotation nodes and multiple columns of standby nodes. Nodes are processed first by the order of levels, and then by groups. The network’s final output is the sum of the bottom nodes’ output of all groups.
The network directs entries of the input vector to different paths for processing and gathers them to form the output. For any given permutation, its network is generated as follows.
Initially, the network contains a single standby node at level 00 and group 00, which stores all entries of the input vector. Each entry is marked with an “unsolved” tag, indicating that it has not yet been rotated to its target position. The network is then constructed group by group. For $i_{g}\geq 0$, the $i_{g}$-th group is constructed as follows.
A group starts from the minimum level containing unsolved entries. For level $\ell\geq 0$, we obtain the maximum remaining rotation distance $rot_{max}$ among all unsolved entries at this level. We set $rot$, the rotation step for the rotation node at level $\ell+1$, to be the largest power of two smaller than $rot_{max}$. Then we arrange each unsolved node’s destination in the next level in 3 steps:
Step 1. We retrieve the entry’s remaining rotation distance $r_{rem}$, original total rotation distance $r_{org}$, and initial position $i$ in the original vector. These values determine the entry’s position in the next level’s node: $i_{dst}=i+(r_{org}-r_{rem})\bmod{n}$.
Step 2. If $r_{rem}\geq rot$, we assign the entry to a rotation node. So we check whether the $i_{dst}$-th position of the rotation node at level $\ell+1$ in the current group is occupied:
  • If not occupied, we let the entry occupy this position. For the edge connecting the entry’s current node and the rotation node, we set its mask’s $i_{dst}$-th position to $1$. The entry’s $r_{rem}$ is updated as $r_{rem}-rot$.
  • Otherwise, we change the entry’s tag from ”unsolved” to ”deferred”, indicating that it will go to level $\ell+1$ in a future group.
Step 3. If $r_{rem}<rot$, we assign the entry to a standby node at level $\ell+1$. Based on the entry’s current node, different decisions are made:
  • If it is a rotation node, we assign the entry to the $m+1$-th standby node where $m$ is the number of standby nodes at level $\ell$. For the edge connecting these two nodes, its mask’s $i_{dst}$-th position is set to $1$.
  • If it is the $k$-th standby node at level $\ell$, the entry is assigned to the $k$-th standby node at level $\ell+1$. For the edge connecting these two nodes, its mask’s $i_{dst}$-th position is set to $1$.
After traversing all unsolved entries, we proceed to the next level by setting $\ell\leftarrow\ell+1$. Otherwise, the $i_{g}$-th group’s construction is completed. We change the tag of entries marked as “unsolved” to “solved” and those marked as “deferred” to “unsolved”, and proceed to the $i_{g}+1$-th group.
Scheme Operation $n$ Number of operations at level $1\rightarrow\log{n}$ $\log{(\#ScalarMult)}$
Ours Rotation $2^{10}$$\{1.0,2.0,3.3,3.7,4.0,4.1,4.3,4.3,4.0,3.8\}$ 31.21
(merged Rescale) $2^{11}$$\{1.0,2.0,3.4,3.9,4.2,4.3,4.3,4.4,4.2,4.1,4.0\}$ 31.39
$2^{12}$$\{1.0,2.0,3.6,4.1,4.2,4.4,4.6,4.8,4.9,4.8,4.7,4.2\}$ 31.53
$2^{13}$$\{1.0,2.0,3.8,4.4,4.7,5.0,5.1,5.2,5.2,5.2,5.1,5.0,4.7\}$ 31.70
$2^{14}$$\{1.0,2.0,4.0,4.7,5.1,5.3,5.4,5.4,5.5,5.6,5.5,5.4,5.3,5.0\}$ 31.83
Benes network Rotation $2^{10}$$\{5.6,6.6,6.6,6.6,6.6,6.6,6.5,6.5,6.5,0\}$ 32.10
(Gentry et al., 2012; Halevi and Shoup, 2014) (double-hoisted) $2^{11}$$\{6.6,3.0,6.6,6.6,6.6,6.6,6.5,6.5,6.5,6.4,0\}$ 32.05
$2^{12}$$\{5.6,6.6,6.6,6.6,6.6,6.6,6.5,6.5,6.5,6.4,6.4,0\}$ 32.25
$2^{13}$$\{6.6,3.0,6.6,6.6,6.6,6.6,6.5,6.5,6.5,6.4,6.4,6.4,0\}$ 32.18
$2^{14}$$\{5.6,6.6,6.6,6.6,6.6,6.6,6.5,6.5,6.5,6.4,6.4,6.4,6.3,0\}$ 32.34
Table 2. Average number of rotations in different network constructions
5.2. Technical Detail
We present the key features and techniques of our proposed network construction.
5.2.1. Multi-group Network Construction
The selection of $rot$ at each level aims to limit both the number of rotations in each group and the total number of rotation keys to at most $logn$. This restriction may cause occupation conflicts when multiple nodes attempt to transfer entries to a single rotation node.
Therefore, new groups are created to manage conflicting entries that can’t secure a position in their original group. Specifically, for any entry encountering an occupation conflict, we track the rotations it has completed and transfer it to one of the subsequent groups using a cross-group edge. Such a strategy ensures that entries follow the paths corresponding to the binary encoding of their total rotation distance.
The number of levels is bounded by $\log{n}+1-i$ for any group in the network, where $i$ is the minimum source level (index) of entries it receives from previous groups. Moreover, the number of levels of groups demonstrates a descending trend since we always consider the closest right-hand group first for transferring conflicting entries.
5.2.2. Reduction of Masks and Depth
Many masks within our network are unnecessary in real homomorphic evaluation. They can be replaced with copy operations without compromising the correctness of the network. For any group with a maximum level index of $\ell$, the unnecessary masks are those on edges reaching standby nodes, except for the edges from level $\ell-2$ to level $\ell-1$. This corresponds to the yellow, blue, and purple edges in Figure 3. Reducing them to copy operations allows us to eliminate a significant number of plaintext-ciphertext multiplications and reduce the multiplication depth of the overall network to at most $\log{n}-1$, which aligns with the requirement of the ideal decomposition.
We further provide a level-collapsing strategy for the network. To compress the top $1\sim\ell$ levels, we apply rotations of distances $\{i\cdot n/2^{\ell}\mid i=0,1,\dots,2^{\ell}-1\}$ to the input ciphertext and assign specific masks to them to construct the inputs for all nodes at level $\ell+1$. The multiplication depth of the merged levels is reduced to $1$. We can also merge the bottom $\ell$ levels by reconstructing the nodes at level $\log{n}-\ell$: for entries sharing the same $r_{rem}$, we use masks to extract and consolidate them into a single ciphertext, yielding $2^{\ell}$ ciphertexts in total. By rotating them according to their respective $r_{rem}$ and summing them, we obtain the result of homomorphic permutation. Rotations in the merged level are performed with a level-order traversal of some $m$-ary tree, where $m$ is a power of $2$. This restricts the number of additional rotation keys required per collapsed level to $\frac{m-1}{\log{m}}-1$.
5.2.3. Timing of Rescale
Rescale is the most computationally intensive operation following rotation in homomorphic permutation, and its execution timing is carefully managed in our proposed network. Whenever a rotation node requires rescaling its input, we reduce its computational complexity by integrating the rescaling into the final submodule of the rotation: ModDown, which receives a ciphertext $ct=(a^{\prime},b^{\prime})\in R_{PQ_{\ell}}^{2}$ from the preceding submodules and outputs $(\lfloor a^{\prime}\cdot P^{-1}\rceil,\lfloor b^{\prime}\cdot P^{-1}\rceil)\in R_{Q_{\ell}}^{2}$ (Cheon et al., 2019; Bossuat et al., 2021). Rescaling a ciphertext $ct=(c_{0},c_{1})\in R_{Q_{\ell}}^{2}$ can be viewed as an instance of ModDown computing $(c_{0},c_{1})\in R_{Q_{\ell-1}}^{2}\leftarrow(\lfloor c_{0}\cdot q_{\ell}^{-1}\rceil,\lfloor c_{1}\cdot q_{\ell}^{-1}\rceil)$. Thus, for a ciphertext requiring both rotation and rescaling, we can substitute the final step of the rotation with a ModDown from $Q_{\ell}P$ to $Q_{\ell-1}$: $(\lfloor a^{\prime}\cdot(q_{\ell}P)^{-1}\rceil,\lfloor b^{\prime}\cdot(q_{\ell}P)^{-1}\rceil)\mod Q_{\ell-1}\in R_{Q_{\ell-1}}^{2}$, completing both rotation and rescaling simultaneously.
$\ell=1$ (Mult Depth=$2$) $\ell=2$ (Mult Depth=$3$) $\ell=3$ (Mult Depth=$4$) $\ell=4$ (Mult Depth=$5$)
Scheme BSGS $d1$ Time (ms) Keys Time (ms) Keys Time (ms) Keys Time (ms) Keys
Original (Jiang et al., 2018) $32$ 1211 38 1108 38 961 38 896 38
Original (Jiang et al., 2018) $16$ 1469 30 1437 30 1202 30 $1$ 30
Ours $16$ 1018 23 722 21 598 21 - -
Ours $8$ 1380 23 844 17 600 15 569 15
Table 3. HMT benchmarks under dimension $d=128$ and different decomposition depth $\ell$
$d=8$ $d=16$ $d=32$ $d=64$
Scheme $d^{\prime}$ $m$ Amortized Keys $m$ Amortized Keys $m$ Amortized Keys $m$ Amortized Keys
Original (Rizomiliotis and Triakosia, 2022) d 64 29 ms 23 8 418 ms 42 1 6142 ms 77 - - -
Ours d 64 17 ms 21 8 184 ms 28 1 1842 ms 35 - - -
Ours d/2 128 13 ms 18 16 127 ms 25 2 1191 ms 32 - - -
Ours d/4 256 7 ms 15 32 86 ms 22 4 870 ms 29 - - -
Ours d/8 512 7 ms 9 64 76 ms 19 8 767 ms 26 1 7594 ms 33
Ours d/16 - - - 128 72 ms 12 16 712 ms 23 2 7009 ms 30
Ours d/32 - - - - - - 32 694 ms 15 4 6762 ms 27
Ours d/64 - - - - - - - - - 8 6829 ms 18
Table 4. HMM benchmarks: full decomposition depth and constant multiplication depth $2$
5.3. Complexity Analysis
In this section, we analyze the primary computational complexity of our proposed multi-group network for homomorphic permutations and compare it with the Benes network-based solution.
As we propose merging Rescale with the ciphertext rotation, the remaining fundamental operations in the network are rotations and homomorphic masking, with the former dominating the complexity. We analyze the complexity of the rotation merged with rescaling, and compare it to the original version of separate routines. For a ciphertext in $R_{Q_{\ell+1}}^{2}$ and modulus chain $P$ with $\alpha$ moduli, the number of scalar multiplication for rescaling it to $R_{Q_{\ell}}^{2}$ and performing one rotation is (see Appendix A for submodules description):
1 (Rescale):
$2\cdot N\cdot((\ell+1)+(\ell+2)\cdot(\log{(N)}+1))$
2 (Decompose):
$N\cdot\log{(N)}\cdot(\ell+1)+\lceil\frac{\ell+1}{\alpha}\rceil\cdot N\cdot(\alpha+(\ell+1)\cdot(\log{(N)}+\alpha))$
3 (MultSum):
$2\cdot N\cdot\lceil\frac{\ell+1}{\alpha}\rceil\cdot(\ell+1+\alpha)$
4 (ModDown):
$2\cdot N\cdot((\ell+1)\cdot(\alpha+\log{(N)}+1)+\alpha\cdot(\log{(N)}+1))$
On the other hand, the number of scalar multiplication for completing the merged rotation is:
1 (Decompose):
$N\cdot\log{(N)}\cdot(\ell+2)+\lceil\frac{\ell+2}{\alpha}\rceil\cdot N\cdot(\alpha+(\ell+2)\cdot(\log{(N)}+\alpha))$
2 (MultSum):
$2\cdot N\cdot\lceil\frac{\ell+2}{\alpha}\rceil\cdot(\ell+2+\alpha)$
3 (ModDown+Rescale):
$2\cdot N\cdot(\ell\cdot((\alpha+1)+\log{(N)}+1)+(\alpha+1)\cdot(\log{(N)}+1)).$
By subtracting the complexity of the merged version from that of the separate version, we obtain a difference no smaller than
$N\cdot(\ell\log N+3.5\log N-\ell+8)$
with $\alpha\geq 2$ under common setting, indicating a complexity reduction across all possible ciphertext modulus levels.
The complexity for computing random permutations of different length $n$ in the network is demonstrated in Table 2, where we present the number of rotations at each level (from $1$ to $\log{n}$). For comparison, the corresponding information for the Benes network-based scheme (Gentry et al., 2012; Halevi and Shoup, 2014) is also provided. In their case, each level corresponds to one double-hoisted homomorphic linear transformation (Bossuat et al., 2021) (see Appendix D for implementation details). The logarithm of the number of scalar multiplications for all rotations in the network is also presented. This is obtained with $logN=15$, modulus chain $Q$ of $18$ moduli, and $P$ of $3$ moduli. At any level $\ell$, the rotations are applied to ciphertexts in $R_{Q_{17-\ell}}^{2}$.
From Table 2, we can observe the following: (i) In our network, the number of rotations first increases and then decreases as the level index grows. The initial increase implies that new groups are created to continue rotating conflicting nodes, while the subsequent decrease arises from the gradual reduction in the number of entries each group needs to process. (ii) For the same $n$, our network exhibits a lower rotation complexity, and this advantage is more pronounced for smaller values of $n$. For $\log{n}=10\sim 14$, the corresponding advantage is approximately $46\%\sim 30\%$.
6. Implementation
We implement the algorithms proposed in the previous sections using the full-RNS CKKS HE scheme (Cheon et al., 2019; Kim and Song, 2019) provided by the Lattigo library (Mouchet et al., 2020). Experiments are conducted on a system with an i7-13700K processor running at 3.40 GHz and a memory of 64G. We present benchmarks for homomorphic matrix transposition, homomorphic matrix multiplication, along with its application in oblivious neural network inference, and arbitrary homomorphic permutations. Unless otherwise specified, the homomorphic encryption parameters are fixed at $\log{QP}\approx 880$ and $\log{N}=15$ across all benchmarks. This provides a 128-bit security level (Albrecht et al., 2015, 2021; Bossuat et al., 2024), an encoding space of $N/2$ complex numbers per ciphertext, and a budget of 18 modulus levels, indexed from 00 to $17$. Note that homomorphic operations at lower modulus levels inherently incur lower computational costs. Thus when we compare different algorithms, we ensure their execution concludes at the same ciphertext modulus level.
6.1. Homomorphic Matrix Transposition
Table 3 provides HMT benchmarks under dimension $d=128$ and decomposition depths $l\leq\log{d}-1$ (requiring multiplication depth $l+1$). In this setup, the leftmost permutation matrix in the decomposition chain of $U^{t}$ has approximately $d/2^{l}$ non-zero diagonals, which is automatically implemented with the BSGS algorithm.
Compared to the original undecomposed version (Jiang et al., 2018), the ideal decomposition yields significant complexity reductions, even when the decomposition depth is shallow. This aligns with the reduction in rotations, whose number changes from $d_{1}+\frac{2d}{d_{1}}$ before decomposition to $2\ell+d_{1}+\frac{2d}{2^{\ell}d_{1}}$ after decomposition. Notably, with $d_{1}=16$, our algorithm demonstrates a speed increase up to 50.2% and a rotation key reduction up to 30.0%.
6.2. Homomorphic Matrix Multiplication
Table 4 compares our enhanced HMM in Section 4.2 to the original scheme (Rizomiliotis and Triakosia, 2022). Unlike our variable $d^{2}d^{\prime}$ matrix encoding space, the original $d^{3}$-encoding allows limited choices of dimensions. Thus we enable comparisons across a broader range of dimensions by using the conjugate-invariant ring based CKKS (Kim and Song, 2019), which shifts the ciphertext’s encoding capacity to $N$ real numbers (HE parameters remain unchanged).
In the table, parameter $m=N/d^{2}d^{\prime}$ denotes the maximum number of matrices that can be computed in parallel. For the original encoding setting $d^{\prime}=d$, our scheme achieves up to a 70% speedup and a 42% reduction in rotation keys compared to the original algorithm. This is attributed to the permutation decomposition and $1$-padding idea, which allow the complexity to grow logarithmically with $d$ and remain independent of multiplication depth overhead. Test results under different $d^{\prime}$ demonstrate the enhancement of our more flexible encoding space utilization method. We can trade encoding space for higher parallelity and lower amortized time per matrix multiplication.
6.3. Application to Oblivious Neural Network Inference
Oblivious inference refers to the task of outsourcing an ML model without revealing its parameters or inputs, wherein the model owner is not required to run the model locally (Jiang et al., 2018; Chen et al., 2019; Rizomiliotis et al., 2022). Jiang et al. (Jiang et al., 2018) proposed an efficient HE-based framework, E2DM, to realize this objective, predicting the MNIST dataset (Deng, 2012) over an encrypted neural network. The computational complexity is dominated by their proposed HMM, which, as discussed above, maintains the best-known amortized complexity. We adopt their service paradigm but replace the HMM module in their framework with our version. Our objective is to demonstrate the practicality of our HMM by achieving lower latency when redundant encoding capacity is available, particularly under small-batch inference scenarios.
Table 5 describes the employed neural network. We call it E2DM-lite, as it is derived by halving the number of convolutional channels and neurons in E2DM, but achieves an accuracy of 97.7% (only 0.4% lower than the original version). We present benchmarks of E2DM-lite in Table 6. When implemented with the original HMM by Jiang et al., E2DM-lite requires an inference batch size of $128$ for optimal performance. Any smaller batch yields no significant time reductions. In contrast, our HMM takes advantage of the encoding space left by small batches, achieving latency reductions of up to 7.9x while maintaining an amortized runtime that is only slightly higher. We provide implementation details in Appendix E.
Table 6 also presents benchmarks of other state-of-art HE-based oblivious neural network inference frameworks on the MNIST dataset. Note that although they use similar underlying HE schemes (Cheon et al., 2017, 2019) and network topologies, we cannot force them to work under identical parameters of HE and neural network, which have been highly customized for different computation strategies. Therefore, the benchmarks are obtained under the original parameters of these frameworks, except for E2DM, which only ensures 80-bit security. We have it implemented under the same HE setting as our scheme.
Jiang et al.’s HMM implementation in E2DM and E2DM-lite exhibits good amortized performance under high batch sizes, but introduces relatively high depth, making it more suitable for shallow, wide networks. The strategy designed by Chen et al.(Chen et al., 2019) (CDKS) employs their proposed multi-key homomorphic encryption to avoid introducing a third-party authority. It has a good latency, but the batch size is relatively small. POSEIDON(Sav et al., 2021) is a federated neural network learning framework based on threshold HE. It provides fast oblivious inference after the network is homomorphically trained among parties. However, clients outside the training system need to encrypt their inquiries with the joint public key of the training parties.
Layer Description
Convolution Input: $28\times 28$, kernel:7×$\times 7$ 7, stride:$3$, channels:$2$
$1^{st}$ square Squaring $128$ inputs
FC-1 Fully connecting $128$ inputs and $32$ outputs
$2^{nd}$ square Squaring $32$ inputs
FC-2 Fully connecting $32$ inputs and $10$ outputs
Table 5. Description of E2DM-lite CNN on the MNIST dataset
Framework Batch Latency Amortized Depth Accuracy
E2DM (Jiang et al., 2018) 64 4232 ms 66 ms 9 98.1%
CDKS (Chen et al., 2019) 1 308 ms 308 ms 6 98.4%
POSEIDON (Sav et al., 2021) 1 155 ms 155 ms 6 89.9%
E2DM-lite* 128 2795 ms 22 ms 9 97.7%
E2DM-lite 32 917 ms 28 ms 6 97.7%
E2DM-lite 16 523 ms 33 ms 6 97.7%
E2DM-lite 8 363 ms 45 ms 6 97.7%
    Original implementation under Jiang et al.’s HMM.
  • Implemented with our HMM.
Table 6. MNIST benchmarks of HE-based ONNI schemes
(Gentry et al., 2012; Halevi and Shoup, 2014) Ours
$\log{n}$ Time (ms) Depth Time (ms) Depth Speed-up
$8$ 1612.21 4 950.74 3 ×$\times 1.69$
$9$ 1816.51 3 1183.75 3 ×$\times 1.53$
$10$ 1687.24 5 1239.73 4 ×$\times 1.36$
$11$ 1909.86 5 1415.83 5 ×$\times 1.35$
$12$ 2198.14 6 1545.78 6 ×$\times 1.42$
$13$ 1887.42 8 1632.65 7 ×$\times 1.15$
$14$ 2278.72 8 1797.25 8 ×$\times 1.27$
Table 7. Average time consumption for randomly sampled permutations of length $n$ ($\log{n}$ rotation keys).
6.4. Arbitrary Homomorphic Permutation
Table 7 compares our proposed homomorphic permutation algorithm in Section 5 with the current optimal approach based on the Benes network (Gentry et al., 2012; Halevi and Shoup, 2014). The average performance on randomly generated permutations with different lengths $n$ is presented in the Table. The benchmarks are constructed as follows: both schemes are initially set with a multiplication depth of $\log{n}-1$ (aligned with the configuration in Table 2). Then, each scheme is applied with its level-collapsing strategy to measure the optimal performance. During the tests, computations with the same $n$ all end at modulus level $117-\log{n}+$ 1.
The Benes network-based solution employs the level-collapsing strategy proposed by Halevi et al.(Halevi and Shoup, 2014). On the other hand, the best practice for our network is to compress the top $2\sim 3$ levels and the bottom $3\sim 4$ levels. Compressing these levels reduces both the multiplication depth and the number of rotations (recall Table 2). Rotations for the compressed levels are further accelerated with the hoisting technique.
The Speed-up column in Table 7 reveals that our algorithm provides speed increase across different $n$. Given the typical setup range of $10\leq\log{N}\leq 16$ for RLWE-based batch-encoding homomorphic encryption schemes (Bossuat et al., 2024), our algorithm demonstrates an efficiency advantage over nearly all practical choices of $n$.
7. Conclusion
In this paper, we employed more effective decompositions on permutations to enhance homomorphic permutation. We proposed two approaches to achieve this: (i) Analyzing and proving the full-depth ideal decomposability of permutations based on our proposed depth-1 ideal decomposition search algorithm. (ii) Designing a new network structure for performing a new form of decomposition and homomorphic computation on arbitrary permutations.
Acknowledgements.
We extend our gratitude to all those who provided valuable insights and suggestions for this paper.
Appendix A Batch-encoding HE Settings
A batch-encoding HE scheme generally has the following setup routines:
  • $\texttt{Setparams}(\cdot)$: Generate parameters $params$: $(N,Q,P,\chi_{sec},\chi_{err})$. $N$ is a power of two, which along with $Q$ and $P$ defines the polynomial rings $R=\mathbb{Z}[X]/(X^{N}+1)$, $R_{Q}=\mathbb{Z}_{Q}[X]/(X^{N}+1)$, and $R_{QP}=\mathbb{Z}_{QP}[X]/(X^{N}+1)$, where $Q=\prod_{i=0}^{L-1}q_{i}$, $P=\prod_{i=0}^{\alpha-1}p_{i}$, and any $q_{i}$ for $0\leq i<L$ or $p_{i}$ for $0\leq i<\alpha$ are congruent to $1$ modulo $2N$. The distributions $\chi_{sec}$ and $\chi_{err}$, defined over $R$, are used for sampling the secret key and noise, respectively. All subsequent routines implicitly receive $params$ as input.
  • $\texttt{KeyPairGen}(\cdot)$: Generate the secret key $s\leftarrow\chi_{sec}$ and compute the public key $(a,-as+e_{1})$, where $a\leftarrow_{U}R_{Q}$.
  • $\texttt{SwitchKeyGen}(s,s^{\prime},\mathbf{w})$ For the input integer decomposition basis of $\beta$ elements, output $swk_{s\leftarrow s^{\prime}}\in R_{Q_{L}P}^{\beta\times 2}$, where $swk_{s\leftarrow s^{\prime}}[i]=(a_{i},-a_{i}s^{\prime}+sw^{(i)}P+e_{i})$ and $a_{i}\leftarrow_{U}R_{Q_{L}P},e_{i}\leftarrow\chi_{err}$.
  • $\texttt{Enc}(m,\ell)$: For the given plaintext $m\in R$ and the index $\ell$ of the modulus chain $Q$, compute $(u,v)\in R_{Q_{\ell}}^{2}\leftarrow(ar+e_{2},(-as+e_{1})r+\Delta_{1}m+\Delta_{2}e_{3})$, where $Q_{\ell}=\prod_{i=0}^{\ell-1}q_{i}$ and $e_{1},e_{2},e_{3}\leftarrow\chi_{err}$. $\Delta_{1}$ and $\Delta_{2}$ are scaling factors used to control the precision of the plaintext, where one of them must be $1$, while the other scales the polynomial coefficients, covering both low-bit and high-bit encoding cases for $m$.
  • $\texttt{Dec}(c,s)$: For the given ciphertext $c=(u,v)\in R_{Q_{\ell}}^{2}$ and secret key $s$, compute $\Delta_{1}m+\Delta_{2}e\leftarrow\left\langle(u,v),(s,1)\right\rangle\in R_{q}$. With appropriately chosen $\Delta_{1}$ and $\Delta_{2}$, one can perform scaling, rounding, and reduction to approximately extract $m$ as the output. For example, setting $\Delta_{1}m=\lfloor\frac{q}{t}m\rfloor$ and $\Delta_{2}=1$ (Fan and Vercauteren, 2012), we obtain $m\in R_{t}\leftarrow\lceil\frac{t}{q}\left\langle(u,v),(s,1)\right\rangle\rfloor\mod{t}$, where $t$ is the plaintext coefficient moduli.
Rotating a ciphertext $ct=(c_{0},c_{1})\in R_{Q_{\ell}}^{2}$ with rotation key $rtk_{k}\in R_{PQ_{L}}^{\beta\times 2}\leftarrow\texttt{SwitchKeyGen}(\phi_{k}(s),s,\mathbf{w})$ requires the following steps (Bossuat et al., 2021), where $\phi_{k}$ is the automorphism for rotating the plaintext vector in $ct$ by $k$ steps:
  1. (1)
    Decompose: $c_{1}$ is switched out of the NTT domain and decomposed based on basis $\mathbf{w}$ of $\beta$ elements. Each decomposed component is then switched back into the NTT domain, forming the vector $\mathbf{d}$.
  2. (2)
    MultSum: The vector $\mathbf{d}$ is inner-multiplied with $rtk_{k}$, yielding the tuple $(a,b)\in R_{PQ_{\ell}}^{2}$.
  3. (3)
    Permute: The tuple $(c_{0}P+a,b)$ undergoes an automorphism $\phi_{k}$, resulting in $(a^{\prime},b^{\prime})\in R_{PQ_{\ell}}^{2}$.
  4. (4)
    ModDown: The computation $(\lfloor a^{\prime}\cdot P^{-1}\rceil,\lfloor b^{\prime}\cdot P^{-1}\rceil)\in R_{Q_{\ell}}^{2}$ is performed.
We use $\beta=\lceil(\ell+1)/\alpha\rceil$ for the complexity analysis in Section 5.3, which is aligned with the RNS decomposition basis in (Bossuat et al., 2021).
Appendix B Section 4 Supplemental Materials
B.1. Partition pattern of arbitary dimension
Theorem 1.
Let $A$ be any square matrix of size $d\times d$ and its initial block size is $d\times d$. Define one round of block partitioning as follows: each block of $A$ with an odd dimension is partitioned and recombined according to equations 11, while each block with an even dimension is partitioned according to equation 5. After any number of rounds of partitioning, $A$ contains blocks of at most two different sizes which differ by only $1$.
Proof.
Suppose that after a certain round of partitioning, the blocks in $A$ have dimensions differing by 1, specifically $d^{\prime}$ and $d^{\prime}+1$. After the next round of partitioning, if $d^{\prime}$ is odd, the blocks of dimension $d^{\prime}$ are divided into blocks of dimensions $\frac{d^{\prime}+1}{2}-1$ and $\frac{d^{\prime}+1}{2}$, while the blocks of dimension $d^{\prime}+1$ are divided into blocks of dimension $\frac{d^{\prime}+1}{2}$. If $d^{\prime}$ is even, the blocks of dimension $d^{\prime}+1$ are divided into blocks of dimensions $\frac{d^{\prime}}{2}+1$ and $\frac{d^{\prime}}{2}$, while the blocks of dimension $d^{\prime}$ are divided into blocks of dimension $\frac{d^{\prime}}{2}$. Given that the two block dimensions initially produced by Equations 11 differ by only 1, by mathematical induction, after each round of partitioning, the blocks in $A$ always have only two different dimensions, differing by only 1. ∎
$d=8$ $d=16$ $d=32$
Scheme $d_{0}$ $m$ Amortized Keys Speed-up $m$ Amortized Keys Speed-up $m$ Amortized Keys Speed-up
Ours (Table 4) $d/1$ 512 7 ms 9 ×$\times 1.00$ 128 72 ms 12 ×$\times 1.00$ 32 694 ms 15 ×$\times 1.00$
Ours (Fast Rep) $d/2$ 512 5 ms 9 ×$\times 1.33$ 128 56 ms 12 ×$\times 1.28$ 32 567 ms 15 ×$\times 1.22$
Ours (Fast Rep) $d/4$ 512 4 ms 12 ×$\times 1.76$ 128 42 ms 15 ×$\times 1.72$ 32 471 ms 18 ×$\times 1.47$
Ours (Fast Rep) $d/8$ 512 2 ms 21 ×$\times 3.13$ 128 30 ms 24 ×$\times 2.35$ 32 354 ms 27 ×$\times 1.96$
Ours (Fast Rep) $d/16$ 128 21 ms 45 ×$\times 3.40$ 32 279 ms 48 ×$\times 2.49$
Ours (Fast Rep) $d/32$ 32 232 ms 93 ×$\times 2.99$
Table 8. Performance of HMM with optimized replication (using $\ell=1,d=d_{1}d_{0}$)
B.2. Demonstration of $1$-padding on $U^{\gamma}$ and $U^{\xi}$
Theorem 2.
For any $d^{2}\times d$ matrices $\hat{A}$ and $\hat{B}$, $U^{\gamma}_{pad}(\hat{A})$ has the same first column as $U^{\gamma}(\hat{A})$, and $U^{\xi}_{pad}(\hat{B})$ has the same first $d$ columns as $U^{\xi}(\hat{B})$.
Proof.
For $U^{\gamma}(\hat{A})$, entries of $A$ in $\hat{A}$ are rotated by $-k(d^{2}-1)$ steps for some $0\leq k<d$ to ensure those with column index $i$ in equivalence class $i\bmod{d}$ are shifted to equivalence class $0\bmod{d}$. Since $(d^{2}-1)$ is coprime to $d$, for each equivalence class $i\bmod{d}$, there is exactly one unique $k$ such that $i+k=0\bmod{d}$. This implies that during each rotation by steps $0,-(d^{2}-1),\dots,-(d^{2}-1)(d-1)$, only one equivalence class is rotated to the first column position of $\hat{A}$. The correctness of $U^{\xi}_{pad}(\hat{B})$ is consistent with the discussion for $U^{\gamma}$ above. $U^{\xi}(\hat{B})$ requires moving $\mathbf{b}_{i}$ in $\hat{B}$ by $k(d-1)$ positions (with $1\times d$ vectors as units) to the column indexed by 0, where $0\leq i,k<d$. For each $\mathbf{b}_{i}$, there is exactly one unique $k$ such that $i+k(d-1)=0\mod d$. This implies that during the rotations of $\hat{B}$ by steps of $0,-(d-1),\dots,-(d-1)^{2}$, each rotation carries exactly one $\mathbf{b}_{i}$ to the first column position. ∎
B.3. Optimization of Replication
In Section 4.2.2, we explore providing HMM with a configurable encoding space of $d^{2}d^{\prime}$. As $d^{\prime}$ decreases, the rotations for replication increases. We propose a method to reduce this complexity. Consider $M=[\mathbf{m}_{0},\mathbf{m}_{1},\ldots,\mathbf{m}_{d-1}]$. To fill the entire $M$ with some $\mathbf{m}_{k}$, a single element replication function $\texttt{SRep}(\mathbf{u}_{k}\odot M;d)$ is typically applied, defined as follows:
$\tilde{M}_{k}=\left[\mathbf{0},\ldots,\mathbf{0},\mathbf{m}_{k},\mathbf{0},\ldots,\mathbf{0},\right]\leftarrow\mathbf{u}_{k}\odot M,$
$\tilde{M}_{k}=\left[\mathbf{m}_{k},\ldots,\mathbf{m}_{k}\right]\leftarrow\texttt{SRep}(\tilde{M}_{k};d):$
$\tilde{M}_{k}\leftarrow\tilde{M}_{k}+\texttt{Rot}(\tilde{M}_{k};i),\text{for}\phantom{a}\log{d}-1\geq i\geq 0.$
This routine requires $O(d\log{d})$ rotations to produce replications of all elements in $M$. To optimize this, we divide $d=d_{1}d_{0}$ and perform rotations with steps not smaller than $d_{0}$ before masking:
(15) $\tilde{M}_{k}\leftarrow\texttt{SRep}\left(\sum^{d_{1}-1}_{i=0}\mathbf{u}_{k,i}\odot\texttt{Rot}\left(M;d_{0}i\right);d_{0}\right),$
where $\mathbf{u}_{k,i}$ extracts $\mathbf{m}_{k}$ from $\texttt{Rot}\left(M;d_{0}i\right)$. The pre-rotated components $\texttt{Rot}(M;d_{0}i)$ can be cached for all $d$ elements’ computation but only costs $d_{1}$ rotations. The total number of rotations is $\left(\frac{d}{d_{0}}+d\log{d_{0}}\right)$, and the multiplication depth remains $1$. For HMM regarding $M$ as a matrix, this method can directly apply to row-wise replication. Extending the pre-rotated steps to {±$\{\pm d_{0}i\mid 0\leq i<d_{1}\}$ enables column-wise replication with a number of $\left(\frac{2d}{d_{0}}+d\log d_{0}\right)$ rotations. Therefore, the number of rotations for our HMM in Section 4.2.2 is reduced to $\left(\frac{3d}{md_{0}}+\frac{2d}{m}\log{d_{0}}\right)$ for $d^{\prime}=1$, surpassing a range of schemes with the complexity of $O(d\log d)$ under $d^{2}$ encoding space (Cheon et al., 2018; Huang and Zong, 2023; Aikata and Roy, 2024). The corresponding performance, along with a comparison to the baseline version without this optimization, is presented in Table 8.
This method can be extended to a multi-layer operation with depth $\ell$. Let $d=d_{\ell}\ldots d_{1}d_{0}$ and $p_{i}=\prod_{j=0}^{i-1}d_{j}$, the equation 15 can be re-written by:
(16) $\tilde{M}_{\lfloor k/p_{\ell}\rfloor}^{(\ell)}\leftarrow\sum^{d_{\ell}-1}_{i=0}\mathbf{u}^{(\ell)}_{\lfloor k/p_{\ell}\rfloor,i}\odot\texttt{Rot}\left(M;p_{\ell}i\right),$
$\tilde{M}_{\lfloor k/p_{j}\rfloor}^{(j)}\leftarrow\sum^{d_{j}-1}_{i=0}\mathbf{u}^{(j)}_{\lfloor k/p_{j}\rfloor,i}\odot\texttt{Rot}\left(\tilde{M}_{\lfloor k/p_{j+1}\rfloor}^{(j+1)};p_{j}i\right),$
$\tilde{M}_{k}^{(1)}\leftarrow\sum^{d_{1}-1}_{i=0}\mathbf{u}^{(1)}_{k,i}\odot\texttt{Rot}\left(\tilde{M}_{\lfloor k/p_{2}\rfloor}^{(2)};p_{1}i\right),$
$\tilde{M}_{k}\leftarrow\texttt{SRep}\left(\tilde{M}_{k}^{(1)};d_{0}\right).$
where $\ell>j>1$. For the $i$-th layer ($0<i$), $d_{\ell}\ldots d_{i+1}(d_{i}-1)$ rotations are required, yielding a rotation complexity of $\left(\frac{d}{d_{0}}+d\log{d_{0}}\right)$ in total. A visualization of the computation process, using $d=4\times 2\times 2=16$ as an example, is shown in Figure 4. Although this extension does not further reduce rotation complexity across depths $\ell$, decomposing $d$ into appropriate factors $d=d_{\ell}\ldots d_{1}d_{0}$ reduces the number of rotation keys and multiplications. For instance, when $d_{0}=1$, the minimum rotational complexity $O(d)$ is achieved, and if $\ell=1$, the rotation keys will be $O(d)$, but the complexity of CMult will increase to $O(d^{2})$. However, if $\ell>1$ and $d_{i}>1$ for $0<i\leq\ell$, the rotation keys are reduced to $O(d_{\ell}+\ldots+d_{2}+d_{1})$. Meanwhile, $d$ ciphertext is generated at the lowest layer ($1$-th layer) and each ciphertext requires $d_{1}$ CMult; $\prod_{j=i}^{\ell}d_{j}$ ciphertext is generated at the $i$-th layer and each ciphertext requires $d_{i}$ CMult; therefore CMult totals $dd_{1}+\sum_{i=2}^{\ell}(d_{i}\prod_{j=i}^{\ell}d_{j})$.
Figure 4. Optimized ciphertext replication by multi-layer construction (using $\ell=2,d=4\times 2\times 2$ as an example)
(a) $16\times 16$ $U^{\sigma}$ with non-zero entries painted in dark blue.
(b) Step 1. Red entries denote the non-zero entries in $U^{\sigma}_{R}$ after Step 1.
(c) Step 2. $U_{R}^{\sigma}$’s final view. Conflicting entries moved away are painted in grey
(d) Step 3. $U^{\sigma}_{L}$ reconstructed by $U^{\sigma}_{R}$ and $U^{\sigma}$.
Figure 5. Applying depth-1 ideal decomposition search to $U^{\sigma}$ of size $16\times 16$.
Figure 6. Demonstration of $U^{\sigma}$ and $U^{\tau}$
Figure 7. Applying $U_{R}^{\sigma}$ and $U_{L}^{\sigma}$ to an $8\times 8$ matrix
Appendix C Full-depth Ideal Decomposability of Permutations in HMM by (Jiang et al., 2018)
Jiang et al. introduced two important permutations in their design of the HMM: $U^{\sigma}$ and $U^{\tau}$. As illustrated in Figures 6, $U^{\sigma}$ transforms the diagonals of the input matrix into columns of the matrix, while $U^{\tau}$ converts the diagonals into rows. Hence, we also refer to $U^{\sigma}$ as the Diag-to-Col transformation and $U^{\tau}$ as the Diag-to-Row transformation. These two permutations account for $5\sqrt{d}$ rotations in the HMM (which has a total of $3d+5\sqrt{d}$ rotations). The full-depth ideal decomposition may reduce this part of complexity to $3\log{d}$ rotations. We also notice that these two permutations have potential applications in bridging homomorphic matrix operations that employ different encoding schemes. Some matrix operation schemes adopt diagonal encoding, which typically maps a single diagonal (using either a single entry or a block as the basic unit) of the matrix into a single ciphertext (Halevi and Shoup, 2014; Huang et al., 2021). If the dataset is initially partitioned and encrypted under row/column-major order, $U^{\sigma}$ and $U^{\tau}$ can be used to reorganize the entries along the diagonal into a row/column, facilitating the extraction of different diagonals into separate ciphertexts.
(a) $16\times 16$ $U^{\tau}$ with non-zero entries painted in dark blue.
(b) Step 1. Red entries denote the non-zero entries in $U^{\tau}_{R}$ after Step 1.
(c) Step 2. No conflicts are generated in the previous step.
(d) Step 3. $U^{\tau}_{L}$ reconstructed by $U^{\tau}_{R}$ and $U^{\tau}$
Figure 8. Applying depth-1 ideal decomposition search to $U^{\tau}$ of size $16\times 16$
C.1. Diag-to-Col Transformation
For any $U^{\sigma}$ with a power-of-2 dimension, the search algorithm finds one of its depth-$1$ ideal decompositions, denoted as $U^{\sigma}=U_{L}^{\sigma}U_{R}^{\sigma}$. Figure 5 shows the decomposition on the $16\times 16$ $U^{\sigma}$. We can deduce the full-depth decomposability from this depth-$1$ solution. For any $d\times d$ matrix $A$ where $d$ is a power of $2$, the effects of applying $U^{\sigma}_{R}$ and $U^{\sigma}_{L}$ to $A$ can be described as follows:
  • $U^{\sigma}_{R}$ evenly partitions $A$ into $4$ blocks of size $d/2\times d/2$ (as indicated in Equation 5), and swaps the lower triangular parts of blocks $A^{\prime}_{0,0}$ and $A^{\prime}_{0,1}$, as well as the upper triangular parts of $A^{\prime}_{1,0}$ and $A^{\prime}_{1,1}$.
  • $U_{L}^{\sigma}$ performs a block-wise diag-to-col transformation on $A$. That is: $U^{\sigma}(A^{\prime}_{0,0}),U^{\sigma}(A^{\prime}_{0,1}),U^{\sigma}(A^{\prime}_{1,0}),U^{\sigma}(A^{\prime}_{1,1})$.
This process, illustrated in Figure 7, can be recursively applied to each block, thus deriving a full-depth ideal decomposition:
(17) $U^{\sigma}(A)=U_{L_{\ell}}^{\sigma}\Big{(}U_{R_{\ell}}^{\sigma}\Big{(}\ldots\Big{(}U_{R_{1}}^{\sigma}(A)\Big{)}\Big{)}\Big{)},$
where $1\leq\ell<\log{d}$. For any $U_{R_{i}}^{\sigma}$ ($0<i<\ell$), it contains only $3$ diagonals, while $U_{L_{\ell}}^{\sigma}$ has only $d/2^{\ell-1}-1$ diagonals (as established by Theorems 1 and 2).
When $A$’s dimension is not a power of 2, the partition method of Equations 10 and 11 is employed to handle blocks with odd dimensions. For the overlapping element $A[d^{\prime},d^{\prime}]$ between $B^{\prime}_{0,0}$ and $B^{\prime}_{1,1}$, $U^{\sigma}(B^{\prime}_{0,0})$ requires it to shift leftward, while $U^{\sigma}(B^{\prime}_{1,1})$ imposes no movement constraints on it. Consequently, this overlap causes no conflicts in element movement. Since each round of decomposition generates at most two distinct block sizes, the corresponding $U_{R}^{\sigma}$ in each round involves no more than $5$ non-zero diagonals.
C.2. Diag-to-Row Transformation
For the $U^{\tau}$ applying to any $d\times d$ matrix $A$, its diagonal distribution is $[0,d^{2}-d]$ with a common difference of $d$. Such a distribution is not symmetric around the zero diagonal. In this case, we may adjust the objective of the search algorithm: let $U_{R}^{\tau}$ only contain non-zero entries at the 00-th and $r_{c}$-th diagonals for $r_{c}=d\cdot\lceil(d-1)/2\rceil$, such that the diagonal distribution of $U_{L}^{\tau}$ converges approximately to $[0,(d^{2}-d)/2]$.
Figure 8 presents the depth-1 decomposition found by the search algorithm. For $d$ as a power of two, the effects of applying $U_{R}^{\tau}$ and $U_{L}^{\tau}$ to $A$ can be described as follows:
  • $U^{\tau}_{R}$ partitions $A$ into left and right halves of size $d\times d/2$: $A=[A^{\prime}_{0},A^{\prime}_{1}]$, and rotates the $A^{\prime}_{1}$ by $\frac{d}{2}\cdot d$ steps, involving only two diagonals.
  • $U_{L}^{\tau}$ applies block-wise diag-to-row transformation to $A$: $U^{\tau}(A^{\prime}_{0})$ and $U^{\tau}(A^{\prime}_{1})$.
By recursively applying the partition and permutations mentioned above to each block, a full-depth ideal decomposition is obtained:
(18) $U^{\tau}(A)=U_{L_{\ell}}^{\tau}\Big{(}U_{R_{\ell}}^{\tau}\Big{(}\ldots\Big{(}U_{R_{1}}^{\tau}(A)\Big{)}\Big{)}\Big{)},$
where $\ell<\log{d}$. For any $U_{R_{i}}^{\tau}$ ($0<i<\ell$), it contains only $2$ diagonals, while $U_{L_{\ell}}^{\tau}$ has only $d/2^{\ell}$ diagonals. When $d$ is not a power of 2, $A$ is divided into $3$ submatrices. For $d=2d^{\prime}+1$, partition $A$ into $A=[A^{\prime}_{0},A^{\prime}_{1},A^{\prime}_{2}]$, where $A^{\prime}_{0}$ and $A^{\prime}_{1}$ are $d\times d^{\prime}$ submatrices, and $A^{\prime}_{2}$ is a $d\times 1$ column vector. $U_{R}^{\tau}$ rotates $A^{\prime}_{1}$ by $d^{\prime}\cdot d$ steps and $A^{\prime}_{2}$ by $2d^{\prime}\cdot d$ steps, resulting in $3$ diagonals. Meanwhile, $U_{L}^{\tau}$ applies the diag-to-row transformation only within the submatrices $A^{\prime}_{0}$ and $A^{\prime}_{1}$: $U^{\tau}(A^{\prime}_{0})$ and $U^{\tau}(A^{\prime}_{1})$.
Appendix D Benes network-based Homomorphic Permutation Implementation
The homomorphic permutation based on the Benes network is implemented with the following details:
  • The Benes network generated for a permutation of length $n$ typically requires a great multiplication depth of $12\log{n}-$ 1 which is unlikely to be practical. Thus, we restrict its depth to $\log{n}-1$ using the optimal level-collapsing method proposed by Halevi and Shoup (Halevi and Shoup, 2014). This depth configuration aligns with our new network construction, which is convenient for comparison. The complexity comparison in Table 2 is obtained with this setting. With depth $\log{n}-1$ being a bottom line, we further search for the best performance of Benes network-based homomorphic permutation by collapsing more levels. This leads to the results in Table 7.
  • The linear transformation at each modulus level is computed using the double-hoisted BSGS algorithm proposed by Bossuat et al. (Bossuat et al., 2021), which improves efficiency by reorganizing and restructuring sub-modules within the rotations. Therefore, the number of rotations in Table 2 is obtained through dividing the number of scalar multiplications for all rotation sub-models in the linear transformation by that of a single rotation.
  • The number of rotation keys for the Benes network generally grows as the levels collapse. Therefore we fix it to roughly $\log{n}$ for a fair comparison with our network. These keys are selected dynamically to minimize the rotation complexity of the Benes network.
HMM by (Jiang et al., 2018) Our HMM
Stage Batch 128 Batch 32 Batch 16 Batch 8
Convolution 121 ms 79 ms 81 ms 82 ms
$1^{st}$ square 31 ms 17 ms 18 ms 16 ms
FC-1 2303 ms 788 ms 379 ms 233 ms
$2^{nd}$ square 15 ms 8 ms 8 ms 8 ms
FC-2 325 ms 25 ms 27 ms 24 ms
Total Latency 2795 ms 917 ms 523 ms 363 ms
Amortized 22 ms 28 ms 33 ms 45 ms
Table 9. Microbenchmarks for E2DM-lite implemented with different HMM schemes
Appendix E Oblivious Neural Network Inference Implementation
We demonstrate how to evaluate E2DM-lite homomorphically. Note that E2DM-lite is designed to be compatible with any HMM that supports parallel computation.
The Convolution Layer. For a kernel $K\in\mathbb{R}^{k\times k}$, a stride of $(s,s)$, $h$ channels, and an input image $I\in\mathbb{R}^{n_{i}\times n_{i}}$, the convolution operation is given by the formula:
$\textbf{Conv}(I,K)_{i^{\prime},j^{\prime}}=\sum_{0\leq i,j<k}K_{i,j}\cdot I_{si^{\prime}+i,\,sj^{\prime}+j}$
for $10\leq i^{\prime},j^{\prime}<d_{K}=\left\lceil\frac{(n_{i}-k)}{s}\right\rceil+$ 1. In E2DM-lite we have $k=7,s=3,h=2,d_{K}=8$, and $n_{i}=28$. The encoding method for input of homomorphic convolution is similar to that of E2DM (Jiang et al., 2018). We first duplicate the set $\{I_{si^{\prime}+i,sj^{\prime}+j}\}_{0\leq i^{\prime},j^{\prime}<d_{K}}$ $h$ times to form one column of a matrix $I^{(i,j)}$, which has $b$ such columns for $b$ images to be processed simultaneously. On the other hand, the set $\{K_{i,j}^{(l)}\}_{0\leq l<h}$ is vertically concatenated and then duplicated $d_{K}^{2}\times b$ times to form a matrix $K^{(i,j)}$ with the same dimensions as $I^{(i,j)}$. Encrypting $I^{(i,j)}$ and $K^{(i,j)}$ into a ciphertext, respectively, for $0\leq i,j<k$, the homomorphic evaluation of the convolution layer can be computed as:
$ct(C)\leftarrow\sum_{0\leq i,j<k}\texttt{Mult}(ct(I^{(i,j)}),ct(K^{(i,j)})),$
where $C\in\mathbb{R}^{hd_{K}^{2}\times b}$ is the output of the layer. Note that the encryption of $I^{(i,j)}$ and $K^{(i,j)}$ follows the encoding rules of the HMM used in the FC-1 layer.
The First Square Layer. This step simply squares the input ciphertext by $ct(C^{2})\leftarrow\texttt{Mult}(ct(C),ct(C))$.
The FC-1 Layer. In the plaintext scenario, this step computes the matrix multiplication of a $w\times hd_{K}^{2}$ weight matrix and an input matrix of size $hd_{K}^{2}\times b$ (for E2DM-lite $w=32$). In the ciphertext scenario, this is achieved through HMM across several submatrices. Specifically, when using an HMM of dimension $d$, we partition $I^{(i,j)}$ (and $K^{(i,j)}$) into $hd_{K}^{2}/d\times b/d$ submatrices, ensuring that $ct(C^{2})$ maintains the same partitioning. We also partition the weight matrix into $w/d\times hd_{K}^{2}/d$ submatrices. This allows us to utilize the parallelism of HMM for $wb/d^{2}$ groups of matrix multiplications, each involving $d\times d$ submatrices of size $hd_{K}/d$. The results are then aggregated in a group-wise order to obtain an output $ct(F)$ with $F$ of size $w\times b$.
The Second Square Layer. This step squares the input ciphertext by $ct(F^{2})\leftarrow\texttt{Mult}(ct(F),ct(F))$.
The FC-2 Layer. In the plaintext scenario, this step computes the matrix multiplication of an $o\times w$ weight matrix and an input matrix of size $w\times b$ (for E2DM-lite $o=10$). Since the partitioning rules are preserved in $ct(F^{2})$, this step is similar to the FC-1 layer, performing $ob/d^{2}$ groups of matrix multiplications, each involving $d\times d$ submatrices of size $w/d$. The results are then aggregated to yield an output $ct(O)$ with $O$ of size $o\times b$. If $d>o$ then zero-padding is applied to the weight matrix.
Table 9 presents benchmarks for E2DM-lite, with $\log{N}=15$ and an encoding space of length $2^{14}$. When using Jiang et al.’s HMM, with $d=32$ and $b=128$, the FC-1 layer performs 4 groups of 4 matrix multiplications of size $32\times 32$, which, with parallelism support, results in a complexity comparable to performing a single $32\times 32$ matrix multiplication. When using our HMM, we set three different parameter sets: $(b=32,d=32,k=4,d_{0}=8)$, $(b=16,d=16,k=4,d_{0}=4)$, and $(b=8,d=8,k=4,d_{0}=2)$. This transforms the FC-1 layer into performing 1, 2, and 4 groups of matrix multiplications of size $32\times 32$, $16\times 16$, and $8\times 8$, respectively. The low latency shown in the table is mainly due to the following reasons: (i) our HMM uses $k=4$-fold redundancy encoding space to provide faster computational speed. (ii) The ciphertext output from the FC-1 layer contains $2^{14}/(32b)$ duplication of the expected matrix of size $32\times b$, which is sufficient to allow the FC-2 layer to collapse into a single multiplication and a recursive summation with $\log{32}$ rotations. This enables the entire E2DM-lite computation to be completed at a lower modulus level, where each layer’s computational cost is lower than that of the HMM implementation by Jiang et al (Jiang et al., 2018).